### v2

# Symmetric Preparation of Systems

Abstract.
In this paper we generalize the Weierstrass and Malgrange preparation theorems to the symmetric matrix valued case, proving symmetric preparation of analytic and smooth symmetric systems that vanishes of first order.
2010 Mathematics Subject Classification:
Primary 32B05 Secondary 47A56, 26B40
1. Introduction and statement of results
If $f$ is $C^{\infty}$ satisfying $\partial_{t}f(0,0)\neq 0$ and $f(0,0)=0$ then by the Malgrange preparation theorem there exist $C^{\infty}$ functions $c(t,x)$ and $r(x)$ such that $c(0,0)\neq 0$, $r(0)=0$ and
(1.1) $f(t,x)=c(t,x)\left(t+r(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$. In the analytic case, the Weierstrass preparation theorem, one has uniqueness but not in the $C^{\infty}$ case, except for real valued $f(x)$ when one can use the implicit function theorem. This also holds for higher order vanishing of $f$ and there are corresponding division results.
The Malgrange preparation theorem has been generalized to $N\times N$ systems of functions in [3]. The simplest example is for smooth $N\times N$ system $F(t,x)$ satisfying $|\partial_{t}F(0,0)|\neq 0$ and $F(0,0)=0$. Then there exist smooth $N\times N$ systems $C(t,x)$ and $M(x)$ such that $|C(0,0)|\neq 0$, $M(0)=0$ and
(1.2) $F(t,x)=C(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$, see [2]. In the analytic case, one has uniqueness but not in the smooth case.
We also proved the following division theorem in [2]. If $F(t,x)$ is a smooth $N\times N$ system, then for any smooth $N\times N$ system $M(x)$ there exist smooth $N\times N$ systems $C(t,x)$ and $R(x)$ such that
(1.3) $F(t,x)=C(t,x)\left(t\operatorname{Id}+M(x)\right)+R(x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
locally near $(0,0)$. As before, one has uniqueness in the analytic case but not in the smooth case. We also have preparation results for higher order vanishing of $F$, adjoint preparation results and corresponding division results, see [3].
Now, the question about symmetric preparation results was raised by Gregory Berkolaiko [1]. The simplest case is for smooth (or analytic) symmetric $N\times N$ matrices $F(t,x)$ satisfying $\partial_{t}F(0,0)>0$ and $F(0,0)=0$. Then, does it exist smooth (or analytic) $N\times N$ matrices $U(t,x)$ and $M(x)$, such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric with $M(0)=0$ satisfying
(1.4) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
near $(0,0)$, and is this preparation unique? Observe that the conditions on $F(t,x)$ are necessary for (1.4) to hold since $U(0,0)U^{*}(0,0)=\partial_{t}F(0,0)>0$ and $F(0,0)=U(0,0)M(0)U(0,0)^{*}=0$. Here and in the following we say that the matrix $A$ is symmetric if $A^{*}=A$ and antisymmetric if $A^{*}=-A$, where $A^{*}=\overline{A}^{t}$.
Remark 1.1.
Observe that the preparation (1.4) is not unique in general. In fact, if $U(t,x)$ and M(x) are solutions to (1.4), then for any orthogonal $A(x)=(A^{-1})^{*}(x)$ we find that $\widetilde{U}(t,x)=U(t,x)A(x)$ and $\widetilde{M}(x)=A^{*}(x)M(x)A(x)$ solves (1.4).
Actually, the preparation is unique in the case when $U$ is symmetric and $U>0$, see Theorem 1.2. We are going to prove the following, which is the main result of the paper.
Theorem 1.2.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ such that
(1.5) $\partial_{t}F(0,0)>0\quad\text{and}\quad F(0,0)=0$
Then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M(x)$ such that $|U|\neq 0$, $M$ is symmetric , $M(0)=0$ and
(1.6) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)\qquad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.6) with analytic $U(t,x)$ and $M(x)$. This preparation is unique if $U$ is symmetric and $U>0$.
Remark 1.3.
Observe that we also get uniqueness in other cases, for example if $U<0$ by taking U-U for for the unique preparations with $U>0$. The proof of Theorem 1.2 also gives the preparation (1.6) for some $0\neq|F(0,0)|\ll 1$ with $|M(0)|\ll 1$, but the bounds depend on the (semi)norms of $F$.
If $F(0,0)\neq 0$ then by subtracting $F(0,0)$ we obtain (1.6) with the remainder $F(0,0)$ giving the following result.
Corollary 1.4.
Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ such that
(1.7) $\partial_{t}F(0,0)>0$
Then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M(x)$ such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric $\forall\,x$, $M(0)=0$ and
(1.8) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)+F(0,0)\qquad\text{in $\omega$}$
If $F(t,x)$ is analytic then we obtain (1.8) with analytic $U(t,x)$ and $M(x)$. This preparation is unique if $U>0$.
Note that the uniqueness only holds for symmetric preparations, since Remark 1.1 gives counterexamples for nonsymmetric $U$. Since (1.6) is not linear in $U(t,x)$ the proof approaches of Weierstrass or Malgrange preparation results do not work in the symmetric case, instead we shall use inverse function theorems. Theorem 1.2 will be proved in Section 3 for the analytic case, and in Section 4 for the smooth case. But first we shall consider the formal power series case.
2. Formal power series solutions
If $F$ is analytic we may solve (1.4) by taking the Taylor expansion of $F(t,x)$:
(2.1) $F(t,x)=\sum_{{j,\alpha}}F_{j,\alpha}t^{j}x^{\alpha}$
where $F_{0,0}=F(0,0)=0$ and $F_{1,0}=\partial_{t}F(0,0)>0$, and take the formal power series expansions $U(t,x)=\sum_{{j,\alpha}}U_{\begin{subarray}{c}j,\alpha\end{subarray}}t^{j}x^{\alpha}$ where $|U_{0,0}|\neq 0$ and $M(x)=\sum_{\alpha\neq 0}M_{\alpha}x^{\alpha}$ is symmetric such that $M_{0}=M(0)=0$. Then (1.4) gives that
(2.2) $\sum_{{j,\alpha}}F_{j,\alpha}t^{j}x^{\alpha}=\sum_{{j,\alpha}}U_{j,\alpha}t^{j}x^{\alpha}\left(t\operatorname{Id}_{N}+\sum_{\gamma\neq 0}M_{\gamma}x^{\gamma}\right)\sum_{{k,\beta}}U_{k,\beta}^{*}t^{k}x^{\beta}$
Simplifying the notations so that $\sum_{{j,\alpha}}A_{\begin{subarray}{c}j,\alpha\end{subarray}}t^{j}x^{\alpha}=\sum_{j,k}A_{j,k}t^{j}x^{k}$, we get
(2.3) $\sum_{j,k}F_{j,k}t^{j}x^{k}=\sum_{j,k}U_{j,k}t^{j}x^{k}\left(t\operatorname{Id}_{N}+\sum_{m>0}M_{m}x^{m}\right)\sum_{{i,\ell}}U_{i,\ell}^{*}t^{i}x^{\ell}$
and by matching coefficients we obtain that
(2.4) $F_{j,k}=\sum_{\begin{subarray}{c}i<j\\ \ell\leq k\end{subarray}}U_{i,\ell}U_{j-1-i,k-\ell}^{*}+\sum_{\begin{subarray}{c}i\leq j\\ \ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{i,\ell}M_{m}U_{j-i,k-m-\ell}^{*}$
The first terms of this expansion are $F_{0,0}=0$, $F_{0,1}=U_{0,0}M_{1}U_{0,0}^{*}$ and $F_{1,0}=U_{0,0}U_{0,0}^{*}$ which gives that $|U_{0,0}|\neq 0$ and $M_{1}=U_{0,0}^{-1}F_{0,1}(U_{0,0}^{-1})^{*}$.
When $j=0$ we obtain from (2.4) that
(2.5) $F_{0,k}=\sum_{\begin{subarray}{c}\ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}\qquad\forall\,k$
so recursion gives
(2.6) $U_{0,0}M_{k}U_{0,0}^{*}=F_{0,k}-\sum_{\begin{subarray}{c}0\leq\ell\leq k-m\\ 0<m<k\end{subarray}}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}\qquad\forall\,k$
We then obtain symmetric $M_{j}$ by
(2.7) $M_{k}=U_{0,0}^{-1}F_{0,k}(U_{0,0}^{-1})^{*}-\sum_{\begin{subarray}{c}0\leq\ell\leq k-m\\ 0<m<k\end{subarray}}U_{0,0}^{-1}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}(U_{0,0}^{-1})^{*}\qquad\forall\,k$
so $M_{k}$ is determined by $F_{0,k}$, $U_{0,\ell}$ with $\ell<k$ and $M_{m}$ with $m<k$.
In general, we obtain from (2.4) with $j$ replaced by $j+1$ that
(2.8) $2\operatorname{Re}U_{j,k}U_{0,0}^{*}=F_{j+1,k}-\sum_{\begin{subarray}{c}0<i<j\\ \ell\leq k\end{subarray}}U_{i,\ell}U_{j-1-i,k-\ell}^{*}-\sum_{\begin{subarray}{c}i\leq j+1\\ \ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{i,\ell}M_{m}U_{j+1-i,k-m-\ell}^{*}$
where the right hand side is determined by $F_{j+1,k}$, $M_{m}$ with $m\leq k$ determined by (2.6) and $U_{i,\ell}$ with $i<j$ and $\ell\leq k$ or $i\leq j+1$ and $\ell\leq k-1$, which gives $i+\ell<j+k$ and $\ell\leq k$ or $i+\ell\leq j+k$ and $\ell\leq k-1$.
Since $|U_{0,0}|\neq 0$, (2.8) can be solved to give $U_{j,k}$ and by Lemma A.1 we get a unique symmetric solution $U_{j,k}$ when $U_{0,0}$ is symmetric. For $k=0$ we find from (2.8) that
(2.9) $2\operatorname{Re}U_{j,0}U_{0,0}^{*}=F_{j+1,0}-\sum_{\begin{subarray}{c}0<i<j\end{subarray}}U_{i,0}U_{j-1-i,0}^{*}$
which gives $2\operatorname{Re}U_{1,0}U_{0,0}^{*}=F_{2,0}$. Since $|U_{0,0}|\neq 0$ we obtain by induction $U_{j,0}$ for any $j$ from (2.9), and if $U_{0,0}$ is symmetric then Lemma A.1 gives a unique symmetric $U_{j,0}$.
Thus, one can determine $U_{j,k}$ and $M_{k}$ from (2.7) and (2.8) by induction over $j+k=p\geq 0$. For $p=0$ we can take any $U_{0,0}$ such that $U_{0,0}U_{0,0}^{*}=F_{1,0}>0$. Then for fixed $p>0$ we get the values of $U_{p,0}$ from (2.9) so we can do a recursion over (pr,r)(p-r,r) for $r=0$ to $p$. If we have determined $U_{j,k}$ for $j+k<p$ and $U_{p-\ell,\ell}$ for $\ell\leq r$ for some $r<p$ then by (2.8) we find that $U_{p-r-1,r+1}$ only depends on $F_{p-r,r+1}$, $U_{i,\ell}$ for $i+\ell\leq p$ and $\ell\leq r$ or $i+\ell<p$ and $\ell\leq r+1$, which are previously determined, and on $M_{m}$ with $m\leq r+1$ which by (2.7) are given by $F_{0,m}$ and $U_{0,\ell}$ with $\ell\leq r<p$ which also have been determined. By recursion over $r$ and induction over $p$ we get the expansion, and if $U_{0,0}$ is symmetric we get unique symmetric $U_{j,k}$ by (2.8) and Lemma A.1, which gives the following result.
Theorem 2.1.
If $F(t,x)$ is analytic and symmetric such that $F(0,0)>0$ and $F(0,0)=0$ then there exists formal power series $U(t,x)$ and $M(x)$ solving (1.4) modulo $O(|t|^{N}+|x|^{N})$, $\forall\,N$. We get a unique power series solution in the case when $U$ is symmetric.
But is it not clear that the formal power expansion converge to an analytic function.
Remark 2.2.
If $n=1$ and $F$ is analytic, then by  [1] the nested Artin approximation (see e.g. Theorem 5.18 in [6]) gives analytic $\overline{U}(x,y)=\sum_{j,k}\overline{U}_{jk}t^{j}x^{k}$ and $\overline{M}(x)=\sum_{k}\overline{M}_{k}x^{k}=\overline{M}^{*}(x)$ that solves
$F(t,x)=\overline{U}(t,x)(tId+\overline{M}(x))\overline{U}^{*}(t,x)\qquad(t,x)\in\mathbf{R}^{2}$
which gives Theorem 1.2 in the two dimensional analytic case. One also obtains that any finite number of terms in the expansions of $\overline{U}(t,x)$ and $\overline{M}(x)$ coincide with the corresponding terms in formal power series solutions $U(t,x)$ and $M(x)$.
When $n>1$ this does not work in general, see the counterexamples in [6]. Instead we will use the inverse function theorem, see Section 3.
3. Proof of Theorem 1.2 in the analytic case
It is clear that $N\times N$ analytic systems form a Banach space with the usual $L^{{}^{\infty}}$ norm, and they can be localized to neighborhoods. We shall prove that if $U(t,x)$ and $M(x)$ are analytic such that $|U|\neq 0$, $M$ is symmetric and $M(0)=0$, then the mapping
(3.1) $(U(t,x),M(x))\overset{\mathcal{F}}{\longrightarrow}\Big(\partial_{t}\big(U(t,x)(t\operatorname{Id}_{N}+M(x))U^{*}(t,x)\big),U_{0}(x)M(x)U_{0}^{*}(x)\Big)\\ =(F_{1}(t,x),F_{0}(x))$
with $U_{0}(x)=U({0,x})$, is analytic and surjective in some neighborhood of $(0,0)$ for analytic and symmetric $F_{j}$, j=0, 1j=0,\,1, when $|F_{1}(0,0)|\neq 0$ and $F_{0}(0)=0$. If we obtain a solution to (3.1) in a neighborhood of $(0,0)$ then we obtain (1.6) with
(3.2) $F(t,x)=F_{0}(x)+\int_{0}^{t}F_{1}(s,x)\,ds\qquad\text{near $(0,0)$}$
thus $F_{0}(x)=F(0,x)$ and $F_{1}(t,x)=\partial_{t}F(t,x)$.
When $M\equiv 0$ we find that $\partial_{t}(tUU^{*})=F_{1}$ so that
(3.3) $U(t,x)U^{*}(t,x)=\frac{1}{t}\int_{0}^{t}F_{1}(s,x)\,ds\to F_{1}(0,x)\text{ when }t\to 0$
For any such solution $U(t,x)$, we find that the differential of $\mathcal{F}$ at $(U,0)$ is given by
(3.4) $(u(t,x),m(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}\big(2t\operatorname{Re}U(t,x)u^{*}(t,x)+U(t,x)m(x)U^{*}(t,x)\big),U_{0}(x)m(x)U_{0}^{*}(x)\Big)\\ =(A_{1}(t,x),A_{0}(x))$
where $u$ and $m$ are analytic and $m=m^{*}$, which gives analytic symmetric $A_{j}$, j=0, 1j=0,\,1. We shall show that this differential is surjective in a neighborhood of $(0,0)$. We find from (3.4) that $m=U_{0}^{-1}A_{0}(U_{0}^{-1})^{*}$ and
$2t\operatorname{Re}U(t,x)u^{*}(t,x)=\int_{0}^{t}A_{1}(s,x)\,ds+R(x)-U(t,x)m(x)U^{*}(t,x)$
for some analytic symmetric $R(x)$. Since the right hand side has to vanish at $t=0$, we take $R(x)=U_{0}(x)m(x)U_{0}^{*}(x)=A_{0}(x)$ and then
$2\operatorname{Re}U(t,x)u^{*}(t,x)=\frac{1}{t}\left(\int_{0}^{t}A_{1}(s,x)\,ds+A_{0}(x)-U(t,x)m(x)U^{*}(t,x)\right)=A(t,x)$
which is analytic and symmetric. By substituting $m$ in $A$ we find that $A$ depends linearly on $A_{j}$, j=0, 1j=0,\,1, and there exists an analytic and antisymmetric $B(t,x)$ so that $2Uu^{*}=A+B$, which gives the solution
(3.5) $u(t,x)=\frac{1}{2}(A(t,x)-B(t,x))(U^{-1})^{*}(t,x)$
Thus (3.4) is surjective so by the inverse function theorem we obtain a local right inverse to $\mathcal{F}$ and local surjectivity of the nonlinear mapping (3.1) in a neighborhood $V$ of $(U,0)$. This neighborhood depend on $U(t,x)$ and thus by $F(t,x)$ by (3.3). Observe that $\mathcal{F}$ will also be surjective on some $(F_{1},F_{0})$ such that $F_{0}(0)\neq 0$.
For any symmetric $M(x)$ such that $M(0)=0$, there exists a neighborhood $\omega$ of $(0,0)$ such that $(U,M)\in V$ at $\omega$, and then (3.5) gives a solution to (3.1) in $\omega$. But (3.4) is not injective, since different antisymmetric $B$ gives different solutions by (3.5). Actually, we have for fixed $U$ and $M\equiv 0$ that the mapping $(A_{0},A_{1},B)\mapsto(u,m)$ is bijective near $(0,0)$.
It remains to prove uniqueness in the case $U(t,x)$ is symmetric and positive definite. Since $m$ already is determined by (3.4), we only need to prove that there is a unique symmetric solution $u(t,x)$ to equation (3.5). In that case, we obtain from Lemma A.1 a unique analytic antisymmetric $B(t,x)$ that solves
$\operatorname{Im}(U^{-1}B)=-\operatorname{Im}(U^{-1}A)\qquad\text{near~$(0,0)$}$
for the analytic $A(t,x)$ in (3.5). Thus the unique analytic symmetric solution is given by
$u(t,x)=\frac{1}{2}U^{-1}(t,x)(A(t,x)+B(t,x))$
We find that $d\mathcal{F}$ is bijective which gives that $\mathcal{F}$ is a analytic bijection near $(0,0)$ in the analytic symmetric case. ∎
4. Proof of Theorem 1.2 in the smooth case
The smooth $N\times N$ systems form a local Fréchet spaces with the $C^{\infty}(\overline{\omega})$ norms, which are tame if $\overline{\omega}$ is a compact manifold with boundary, see Theorem C.8. For tame Fréchet spaces one can use the Nash–Moser theorem, which is the inverse function theorem for tame Fréchet spaces, see Theorem C.12. The difference to the analytic case is that is it not enought to check the local surjectivity of the linearized map when $M\equiv 0$, instead one has to check that for any $U$ and $M$ in a neighborhood.
As before, it suffices to show that if $|U(0,0)|\neq 0$, $M(x)$ is symmetric and $M(0)=0$ then the following $C^{\infty}$ mapping is surjective near $(0,0)$
(4.1) $C^{\infty}\ni(U(t,x),M(x))\overset{\mathcal{F}}{\longrightarrow}\\ \left(\partial_{t}\big(U(t,x)(t\operatorname{Id}_{N}+M(x))U^{*}(t,x)\big),U_{0}(x)M(x)U_{0}^{*}(x)\right)\\ =(F_{1}(t,x),F_{0}(x))\in C^{\infty}$
where $U_{0}(x)=U({0,x})$. Here $F_{0}$ and $F_{1}$ are symmetric, $F_{1}(0,0)>0$ and $F_{0}(0)=0$. By (3.2) we then obtain $F$ from $F_{0}$ and $F_{1}$. It follows from Example C.4 that the mapping $\mathcal{F}$ is tame of degree 1.
For any such $(F_{1},F_{0})$ we may take smooth $(U(t,x),M(x))$ so that $U(0,0)U^{*}(0,0)=F_{1}(0,0)>0$ and $M(0)=0$. Since we only need local solutions, we may assume that $U(t,x)\in C^{\infty}(\overline{\omega}_{n+1})$ so that $U(0,0)$ is invertible, $M(x)\in C^{\infty}(\overline{\omega}_{n})$ is symmetric and $M(0)=0$, where $\overline{\omega}_{k}\subset\mathbf{R}^{k}$ is a compact neighborhood of the origin.
Let $V$ be a bounded neighborhood of $(U(t,x),M(x))$ in $C^{\infty}(\overline{\omega}_{n+1})\times C^{\infty}(\overline{\omega}_{n})$ so that the first component is invertible and the second component is symmetric. Since $M(0)=0$ we may for any $\varepsilon>0$ choose $V$ and a compact neighborhood $\overline{\omega}$ of the origin so that $(\overline{U}(t,x),\overline{M}(x))\in V$ gives $\|\overline{M}(x)\|\leq\varepsilon$ when $x\in\overline{\omega}$. For any $(U(t,x),M(x))\in V$ the differential of $\mathcal{F}$ at $(U(t,x),M(x))$ is equal to
(4.2) $C^{\infty}\ni(u(t,x),m(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \left(\partial_{t}\big(2\operatorname{Re}(U(t\operatorname{Id}_{N}+M(x))u^{*})+UmU^{*}\big),U_{0}mU_{0}^{*}+2\operatorname{Re}(U_{0}Mu_{0}^{*})\right)\\ =(A_{1}(t,x),A_{0}(x))\in C^{\infty}$
where $u_{0}(x)=u(0,x)$, and $m$, $A_{0}$ and $A_{1}$ are symmetric. We shall show that if $\overline{\omega}_{n+1}$ and $\overline{\omega}_{n}$ are sufficiently small, then the mapping $d\mathcal{F}$ is tame and surjective with a tame right inverse, which will show the surjectivity of $\mathcal{F}$ by Theorem C.12. For that, we need that $\|M(x)\|\ll 1$ for $|x|\ll 1$.
We find from (4.2) that
(4.3) $m=U_{0}^{-1}\Big(A_{0}-2\operatorname{Re}(U_{0}Mu^{*}_{0})\Big)(U_{0}^{-1})^{*}$
is linear in $Mu_{0}^{*}$ and if $P(t,M(x))=t\operatorname{Id}_{N}+M(x)$ then
(4.4) $2\operatorname{Re}UP(t,M(x))u^{*}=\int_{0}^{t}A_{1}(s,x)\,ds-U(t,x)m(x)U^{*}(t,x)+R(x)=A(t,x)$
for some smooth symmetric $R(x)$. Observe that $A$ depends linearly on $m$ and thus linearly on $Mu_{0}^{*}$ by (4.3). By (4.4) there exists a smooth antisymmetric $B(t,x)$ so that
(4.5) $2P(t,M(x))u^{*}(t,x)=\\ U^{-1}(t,x)\int_{0}^{t}A_{1}(s,x)\,ds-m(x)U^{*}(t,x)+U^{-1}(t,x)R(x)+U^{-1}(t,x)B(t,x)$
Since $R(x)$ is symmetric and $B(0,x)$ is antisymmetric, we may replace $R(x)$ with $R(x)+B(0,x)$ and let $B(0,x)\equiv 0$. Then we can replace $B(t,x)$ with $tB(t,x)$, which gives
(4.6) $U^{-1}(t,x)tB(t,x)=P(t,M(x))U^{-1}(t,x)B(t,x)-M(x)U^{-1}(t,x)B(t,x)$
Now we shall use the symmetric Malgrange division theorem given by Corollary B.4 on every term on the right hand side of (4.5) using (4.6). That will give a $C^{\infty}$ surjective tame map with a $C^{\infty}$ tame right inverse by Example C.13.
First, if $P=P(t,M(x))$ then Corollary B.4 gives smooth $C_{1}(t,x)$ and $R_{1}(x)$ so that
(4.7) $U^{-1}(t,x)\int_{0}^{t}A_{1}(s,x)\,ds=PC_{1}(t,x)+R_{1}(x)\qquad\text{near $(0,0)$}$
where $C_{1}$ and $R_{1}$ depend linearly on $m$ and thus linearly on $Mu_{0}^{*}$ by (4.3), and that
(4.8) $m(x)U^{*}(t,x)=PC_{2}(t,x)+R_{2}(x)\qquad\text{near $(0,0)$}$
where $C_{2}$ and $R_{2}$ are smooth and depend linearly on $Mu_{0}^{*}$. We also find
(4.9) $U^{-1}(t,x)=PC_{3}(t,x)+R_{3}(x)\qquad\text{near $(0,0)$}$
with smooth $C_{3}$ and $R_{3}$ such that $R_{3}(0)=U^{-1}(0,0)$ is invertible, thus $R_{3}(x)$ is invertible when $|x|\ll 1$. Finally, we obtain smooth $C_{4}(t,x)$ and $R_{4}(x)$ so that
(4.10) $M(x)U^{-1}(t,x)B(t,x)=PC_{4}(t,x)+R_{4}(x)\qquad\text{near $(0,0)$}$
and $C_{4}(t,0)=R_{4}(0)=0$. Here $C_{4}$ and $R_{4}$ depend linearly on $B$ but the differentials with respect to $B$ is $\mathcal{O}(M)$. We have that $C_{j}$ and $R_{j}$ depend tamely on $U$ and $M$, $\forall\,j$.
Summing up, we find from (4.6)–(4.10) that
(4.11) $2Pu^{*}(t,x)=P(C_{1}+C_{2}+C_{3}R+C_{4}+U^{-1}B)+R_{1}+R_{2}+R_{3}R+R_{4}=PC_{0}+R_{0}$
near $(0,0)$, where $C_{0}(t,x)$ and $R_{0}(x)$ are smooth depending linearly on $R$ $Mu_{0}^{*}$ and $B$. Thus for fixed $U$, $M$ and $B$, we can write
(4.12) $\left\{\begin{aligned} &C_{0}=C_{00}+C_{01}Mu_{0}^{*}+C_{02}R+C_{03}B+U^{-1}B\\ &R_{0}=R_{00}+R_{01}Mu_{0}^{*}+R_{3}R+R_{03}B\end{aligned}\right$.
where $R_{3}$ and $U^{-1}$ are invertible. Observe that, except $U^{-1}B$ and $R_{3}R$, the products are not matrix multiplications, instead they represent linear maps. We have that the norms $\|C_{03}\|$ and $\|R_{03}\|$ are bounded by $C\|M\|$.
Now, since $u_{0}(x)$ appears in the equation we shall let $G(t,x)=G_{0}(x)+tG_{1}(t,x)$ for any smooth $G$, and $C_{jk}(t,x)=C_{jk}^{0}(x)+tC_{jk}^{1}(t,x)$, $\forall\,jk$. For fixed $U$, $M$ and $B$ we obtain that $2u^{*}\equiv C_{0}$ with $R_{0}\equiv 0$ if
(4.13) $\left\{\begin{aligned} &2u_{0}^{*}-C_{01}^{0}Mu_{0}^{*}-C_{02}^{0}R\equiv C_{00}^{0}+(C_{03}B)_{0}+(U^{-1}B)_{0}\\ &-tC_{01}^{1}Mu_{0}^{*}+2tu_{1}^{*}-tC_{02}^{1}R\equiv tC_{00}^{1}+t(C_{03}B)_{1}+t(U^{-1}B)_{1}\\ &R_{01}Mu_{0}^{*}+R_{3}R\equiv-R_{00}-R_{03}B\end{aligned}\right$.
By dividing the second equation by $t$, this system has the following matrix for $(u^{*}_{0},u_{1}^{*},R)$
(4.14) $\begin{pmatrix}2\operatorname{Id}_{N}-C_{01}^{0}M&0&-C_{02}^{0}\\ -C_{01}^{1}M&2\operatorname{Id}_{N}&-C_{02}^{1}\\ R_{01}M&0&R_{3}\end{pmatrix}$
which is invertible when $M\equiv 0$ since $R_{3}$ is invertible, and therefore when $\|M(x)\|\ll 1$. Thus, if the neighborhood $V$ is small enough, then for any $B$ we obtain a unique solution $u^{*}\equiv u_{0}^{*}+tu_{1}^{*}\equiv C_{0}/2$ with $R_{0}\equiv 0$ in a sufficiently small compact neighborhood $\overline{\omega}$ of $(0,0)$. This gives a smooth tame map by Example C.4, and the neighborhood $\overline{\omega}$ only depends on the bounds on $U$, $M$ and $B$.
Having found $u^{*}$ we obtain $(u,m)$ by (4.3) in $\overline{\omega}$. This gives surjectivity and a tame right inverse to $d\mathcal{F}$ locally for any $(U,M)\in V$. By Theorem C.12, we obtain that the $C^{\infty}$ mapping $\mathcal{F}$ has a tame local right inverse on a neighborhood of $(F_{1},F_{0})$ in (4.1), giving a solution to (1.6). Observe that $\mathcal{F}$ is surjective also on some $(F_{1},F_{0})$ such that $F_{0}(0)\neq 0$ depending on $F_{1}$. Note that every antisymmetric $B$ gives a solution.
It remains to prove uniqueness in the symmetric case for $(U,M)\in V$ when $U(t,x)>0$ and $u(t,x)$ is symmetric. We shall then show that there is only one symmetric solution to (1.6), by finding the corresponding $B(t,x)$ in (4.5). By adding the equation $\operatorname{Im}u^{*}\equiv\operatorname{Im}(C_{0}/2)\equiv 0$ with $C_{0}$ given by (4.12), we find
(4.15) $\left\{\begin{aligned} &2u_{0}^{*}-C_{01}^{0}Mu_{0}^{*}-C_{02}^{0}R-(C_{03}B)_{0}-(U^{-1}B)_{0}\equiv C_{00}^{0}\\ &-tC_{01}^{1}Mu_{0}^{*}+2tu_{1}^{*}-tC_{02}^{1}R-t(C_{03}B)_{1}-t(U^{-1}B)_{1}\equiv tC_{00}^{1}\\ &R_{01}Mu^{*}+R_{3}R+R_{03}B\equiv-R_{00}\\ &\operatorname{Im}(C_{01}Mu_{0}^{*})+\operatorname{Im}(C_{02}R)+\operatorname{Im}(C_{03}B)+\operatorname{Im}(U^{-1}B)\equiv-\operatorname{Im}(C_{00})\end{aligned}\right$.
By dividing the second equation by $t$, this system has following matrix for $(u_{0}^{*},u_{1}^{*},R,B)$
(4.16) $\begin{pmatrix}2\operatorname{Id}_{N}-C_{01}^{0}M&0&-C_{02}^{0}&-(C_{03}\,\boldsymbol{\cdot})_{0}-(U^{-1}\,\boldsymbol{\cdot})_{0}\\ -C_{01}^{1}M&2\operatorname{Id}_{N}&-C_{02}^{1}&-(C_{03}\,\boldsymbol{\cdot})_{1}-(U^{-1}\,\boldsymbol{\cdot})_{1}\\ R_{01}M&0&R_{3}&R_{03}\\ \operatorname{Im}(C_{01}M\,\boldsymbol{\cdot})&0&\operatorname{Im}(C_{02}\,\boldsymbol{\cdot})&\operatorname{Im}(C_{03}\,\boldsymbol{\cdot})+\operatorname{Im}(U^{-1}\,\boldsymbol{\cdot})\end{pmatrix}$
where $C\,\boldsymbol{\cdot}$ is the map $E\to CE$. Thus $\operatorname{Im}(U^{-1}\,\boldsymbol{\cdot})$ is the map $E\mapsto\operatorname{Im}(U^{-1}E)$, which is invertible by Lemma A.1 because $U^{-1}>0$. Since $R_{3}$ also is invertible and $C_{03}=R_{03}=0$ when $M\equiv 0$, we find that (4.16) is invertible when $M\equiv 0$, and therefore when $\|M(x)\|\ll 1$. Thus, if the neighborhood $V$ is small enough, then we obtain a unique smooth solution $(u_{0}^{*},u_{1}^{*},R,B)$ to (4.15) near $(0,0)$ giving symmetric $u\equiv u_{0}+tu_{1}\equiv C_{0}/2$ with $R_{0}\equiv 0$. By (4.3) this gives a unique symmetric solution $(u,m)\in C^{\infty}$ to (4.2). Observe that this gives a tame map by Example C.4, and the neighborhood $\overline{\omega}$ only depends on the bounds on $(U,M)\in V$. Thus we obtain in the positive definite case that $d\mathcal{F}$ is a tame bijection locally for any positive definite $U$ and symmetric $M$ such that $(U,M)\in V$, so by Theorem C.12 we obtain that the mapping $\mathcal{F}$ is a tame bijection locally in this case, giving a unique solution to (1.6). ∎
Appendix A A symmetry lemma
In this section, we are going to prove the following result about symmetric and antisymmetric invariance of matrices.
Lemma A.1.
If $U(x)>0$ is a smooth $N\times N$ matrix, then $A(x)\mapsto\operatorname{Re}(UA)(x)$ is a bijection of smooth symmetric matrices and $A(x)\mapsto\operatorname{Im}(UA)(x)$ is a bijection of smooth antisymmetric matrices. If $U(x)$ is analytic, then this holds for analytic $A(x)$.
Proof.
First we consider the case of constant matrices. To simplify, we define the symmetric product $S(U,A)=\frac{1}{2}(UA+AU)$, which is symmetric when $A$ is symmetric and antisymmetric when $A$ is antisymmetric, thus $A\mapsto S(U,A)$ preserves the symmetry. Since any $N\times N$ matrix can be written uniquely as a sum of a symmetric and an antisymmetric matrix, we find that $A\to S(U,A)$ is a bijection for any $N\times N$ matrix $A$. We also have that linear ON mappings $O$ preserves the symmetry, since $OAO^{-1}=OAO^{*}$.
Thus, by using an ON base of eigenvectors to $U$, we find that $U=\{\mu_{j}\delta_{jk}\}_{jk}$ with $\mu_{j}>0,\ \forall\,j$, and $A=\{\alpha_{jk}\}_{jk}$. If $A$ is symmetric then $\alpha_{kj}=\overline{\alpha}_{jk}$ $\forall\,jk$, and if $A$ is antisymmetric then $\alpha_{kj}=-\overline{\alpha}_{jk}$ $\forall\,jk$. In both cases, we find that
$S(U,A)=\{(\mu_{j}+\mu_{k})\alpha_{jk}/2\}_{jk}=0$
if and only if $\alpha_{jk}=0$ $\forall\,jk$. So $A\mapsto S(U,A)$ is injective, and thus bijective in both cases. Observe that $|\alpha_{jk}|\leq\|U^{-1}\|\|S(U,A)\|$ since $\mu_{j}^{-1}\leq\|U^{-1}\|$, $\forall\,j$.
Next, we consider the equation
$S(U,A)(x)=S(U(x),A(x))=B(x)\in C^{\infty}$
which has a unique solution $A(x)=\{\alpha_{jk}\}_{jk}(x)$ such that $|\alpha_{jk}(x)|\leq\|U^{-1}(x)\|\|B(x)\|$. By compactness we find for any $x_{0}$ and any sequence $x_{j}\to x_{0}$ that there exixts a subsequence having a limit: $A(x_{j})\to A_{0}$. Then by continuity and uniqueness we find that $\lim_{x\to x_{0}}A(x)=A_{0}=A(x_{0})$, which proves continuity of $A(x)$.
To show that $A(x)$ is differentiable, we put $\Delta_{h,y}A(x)=\frac{A(x+hy)-A(x)}{h}$ for any $x,\,y\in\mathbf{R}^{n}$ and $h>0$. Then we find
(A.1) $\Delta_{h,y}S(U,A)(x)=S(\Delta_{h,y}U(x),A(x+hy))+S(U(x),\Delta_{h,y}A(x))=\Delta_{h,y}B(x)$
which gives that
(A.2) $S(U(x),\Delta_{h,y}A(x))=\Delta_{h,y}B(x)-S(\Delta_{h,y}U(x),A(x+hy))$
where the right hand side is continuous in $h$. As before, we that the left hand side has the unique limit $S(U(x_{0}),y\cdot\partial A(x_{0}))$ for any $y$, giving $S(U,\partial A)=\partial B-S(\partial U,A)$.
Let $A_{\alpha}(x)=\partial^{\alpha}A(x)$ and $A_{k}(x)=\{A_{\alpha}(x)\}_{|\alpha|\leq k}$ and similarly for $U(x)$ and $B(x)$. By induction, we can assume that $A(x)\in C^{k}$ and that
(A.3) $B_{\gamma}(x)=\sum_{\alpha+\beta=\gamma}c^{\gamma}_{\alpha\beta}S(U_{\alpha}(x),A_{\beta}(x))\qquad\text{for $|\gamma|=k$}$
Then we find for $|\beta|=k$ that
(A.4) $S(U(x),\Delta_{h,y}A_{\beta}(x))=\Delta_{h,y}B_{\beta}(x)-S(\Delta_{h,y}U(x),A_{\beta}(x+hy))\\ -\Delta_{h,y}\sum_{\alpha\neq 0}c_{\alpha\beta}S(U_{\alpha}(x),A_{\beta-\alpha}(x))$
where the right hand side is continuous in $h$. As before, we find that that the left hand side has the limit $S(U,y\cdot\partial A_{k})(x)$ which gives that $A_{k}(x)\in C^{1}$ so $A(x)\in C^{k+1}$. By induction we obtain the result in the $C^{\infty}$ case.
If $U(x)$ and $B(x)$ are analytic, then they have holomorphic extensions in a complex neighborhood of $\mathbf{R}^{n}$. We then find that $S(U(z),\partial_{\bar{z}_{j}}A(z))=\partial_{\bar{z}_{j}}B(z)-S(\partial_{\bar{z}_{j}}U(z),A(z))=0$ in that neighborhood $\forall\,j$, which gives the analyticity of $A(x)$ by uniqueness. ∎
Appendix B Symmetric Malgrange division
In this section, we shall prove a symmetric version of the Malgrange preparation theorem for systems, by modifying the proofs in the Appendix of [2] which generalized [5, Sect. 7.5] to systems. In what follows, we shall let $|A|=\det A$ be the determinant of $N\times N$ matrices $A$, and let $\|A\|$ be the matrix norm.
We are going to divide matrix valued analytic functions with symmetric matrix valued (first order) polynomials. We shall let $P(t,B)=t\operatorname{Id}_{N}+B$, where $B$ is symmetric $N\times N$ matrix, $\omega$} be an open set in $\mathbf{C}$ such that $|P(t,B)|\neq 0$ when $t\in\partial{\omega}\in C^{1}$. If $G(t)$ is analytic $N\times N$ system in $\overline{\omega}$, then we have
(B.1) $G(t)=Q(t)P(t,B)+R{\qquad}t\in{\omega}$
where
(B.2) $Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds{\qquad}t\in{\omega}$
is analytic and
(B.3) $R$ $=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(P(s,B)-P(t,B))(s-t)^{-1}\,ds$
$=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}\,ds$
is constant. Observe that $Q(t)$ and $R$ depend analytically on $B$. The remainder RR\/ is uniquely determined by the condition that $RP(t,B)^{-1}$ is analytic when $t\not\in{\omega}$. In fact, it is easy to see that (B.3) gives
$RP(t,B)^{-1}=-(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds{\qquad}t\not\in{\omega}$
which is analytic in $\complement{\omega}$. Then the uniqueness of RR\/ follows, since (B.1) implies that
$Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds-(2{\pi}i)^{-1}\int_{\partial{\omega}}RP(s,B)^{-1}(s-t)^{-1}\,ds$
when $t\in{\omega}$. If $RP(s,B)^{-1}$ is analytic in $\complement{\omega}$ we may push the integration contour to infinity in the second integral. Since the integrand is $O(|s|^{-2})$ in the first integral we obtain (B.2). Now we want to make this division in a small complex strip around $\mathbf{R}$, with uniform bounds.
Lemma B.1.
Let $G(t)$ be a bounded analytic $N\times N$ system in $\{\,t\in\mathbf{C}:|\operatorname{Im}t|<{\varepsilon}\,\}$. Then for any symmetric $N\times N$ matrix $B$ we can find analytic $N\times N$ system $Q^{{\varepsilon}}(t,B,G)$ and $R^{{\varepsilon}}(B,G)$ depending linearly on $G(t)$, such that
(B.4) $G(t)=Q^{{\varepsilon}}(t,B,G)P(t,B)+R^{{\varepsilon}}(B,G){\quad}\text{if}{\quad}\|B\|<1$
We also have the following estimates
(B.5) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q^{{\varepsilon}}(t,B,G)\|\leq C_{{\alpha}{\beta}}M{\varepsilon}^{-{\alpha}-|{\beta}|-2}$
$\|\partial_{B}^{{\beta}}R^{{\varepsilon}}(B,G)\|\leq C_{{\beta}}M{\varepsilon}^{-|{\beta}|-1}$
where $M=\sup\{\,\|G(s)\|:|\operatorname{Im}s|\leq{\varepsilon}\,\}$ and the constants are independent of ${0<\varepsilon}\leq 1$.
Observe that for the general case when $B$ is not symmetric, we find that the estimates (4.5) have the factor $N$ in front of $|\beta|$, see [2, Lemma A.1].
Proof.
This will be an adaptation of the proof of Lemma A.1 in [2]. Let $A$ be symmetric $N\times N$ matrix, and assume $\|A\|<1$. This gives that 1<t<1-1<t<1 when $|P(t,A)|=0$ since then tw=Awtw=-Aw for some $w\neq 0$. Thus we have that
(B.6) $|P(t,A)|\neq 0\qquad\text{when $\operatorname{Im}t\neq 0$}$
Let $\omega$} be the rectangle $\{\,s:|\operatorname{Im}s|<{\varepsilon},|\operatorname{Re}s|<2\,\}$, $0<\varepsilon\leq 1$, and let ${\omega}_{t}$ be the union of this rectangle and a congruent one centered at $\operatorname{Re}t$. Then (B.1) is satisfied when $|\operatorname{Im}t|<{\varepsilon}$, with
(B.7) $Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}_{t}}G(s)P(s,A)^{-1}(s-t)^{-1}\,ds$
and
(B.8) $R=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,A)^{-1}\,ds$
Since $P(t,A)$ is symmetric, we find from (B.6)
(B.9) $\|P(s,A)^{-1}\|\leq{\varepsilon}^{-1}\quad\text{on}\quad\partial{\omega}_{t}\bigcup\partial{\omega}$
In fact, if $\alpha_{j}\in(-1,1)$ are the eigenvalues of $A$ then the eigenvalues of $P(t,A)$ are $t+\alpha_{j}$, and $\|P(t,A)^{-1}\|\leq\max_{j}|t+\alpha_{j}|^{-1}\leq\varepsilon^{-1}$ when $|\operatorname{Im}t|\geq\varepsilon$ or $|\operatorname{Re}t|\geq 2$. This gives the estimates
(B.10) $\left\{\begin{aligned} &\|Q(t)\|\leq CM{\varepsilon}^{-2}\quad\text{when $|\operatorname{Im}t|<{\varepsilon}$}\\ &\|R\|\leq CM{\varepsilon}^{-1}\end{aligned}\right$.
Now we shall perturb the coefficients of $P$. Since $\|P(s,A)^{-1}\|\leq{\varepsilon}^{-1}$ on $\partial{\omega}\cup\partial{\omega}_{t}$, we find $\|P(s,B)^{-1}\|\leq c{\varepsilon}^{-1}$ there, if $\|B-A\|\leq c_{0}{\varepsilon}$ for some $c_{0}\ll 1$. In fact, the eigenvalues of symmetric matrices depends Lipschitz continuously on the matrix norms by the minimax principle. Thus, by using (B.7)–(B.8) for P(t,B)P(t,B)\/ we obtain (B.4) with $Q^{{\varepsilon}}=Q$ and $R^{{\varepsilon}}=R$ when $\|B-A\|\leq c_{0}{\varepsilon}$. Clearly, QQ\/ and RR\/ are analytic functions of $B$. We easily obtain the estimates (B.5) when $|\operatorname{Im}t|<{\varepsilon}$ and $\|B-A\|\leq c_{0}{\varepsilon}$.
Next, take a suitable partition of unity $\{\,{\Psi}_{j}(X)\,\}$ for the symmetric matrices, which can be parametrized by $\mathbf{R}^{N}\times\mathbf{C}^{N(N-1)/2}\cong\mathbf{R}^{N^{2}}$, for example the one in [5, Theorem 1.4.6], and put ${\Phi}_{j}(X)={\Psi}_{j}(X/c{\varepsilon})$. If $c$ is small enough (independently of $\varepsilon$}), we obtain $\|X-Y\|<c_{0}{\varepsilon}$ if $X$, $Y\in\operatorname{\rm supp}{\Phi}_{j}$, $\forall\,j$. By doing the above construction with $A_{j}\in\operatorname{\rm supp}{\Phi}_{j}$, we get the functions $Q^{{\varepsilon}}_{j}$ and $R^{{\varepsilon}}_{j}$. By taking $Q^{{\varepsilon}}=\sum_{j}{\Phi}_{j}Q^{{\varepsilon}}_{j}$ and $R^{{\varepsilon}}=\sum_{j}{\Phi}_{j}R^{{\varepsilon}}_{j}$, we obtain the result. ∎
The proof of Proposition A.2 in [2] now gives the following result.
Proposition B.2.
Let $G(t)\in\mathcal{S}(\mathbf{R})$ with values in $N\times N$ matrices and let $B$ be a symmetric $N\times N$ matrix. Then we can find smooth $N\times N$ systems $Q(t,B,G)$ and $R(B,G)$ depending linearly on $G(t)$, such that (B.4) holds with QQ\/ and RR\/ instead of $Q^{{\varepsilon}}$ and $R^{{\varepsilon}}$. Also, we have the following estimates
(B.11) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q(t,B,G)\|\leq C_{{\alpha}{\beta}}\int(\|G\|+\|G^{(\nu)}\|)\,dt{\quad}\nu={\alpha}+|{\beta}|+4$
$\|\partial_{B}^{{\beta}}R(B,G)\|\leq C_{{\beta}}\int(\|G\|+\|G^{(\nu)}\|)\,dt{\quad}\nu=|{\beta}|+3$
Observe that in the general case, when $B$ is not symmetric, the number of derivatives of $G$ in (B.11) grows as $N|\beta|$ so the estimates are not tame in general, see Appendix C. For completeness, we shall prove Proposition B.2.
Proof.
Choose a function $\psi\in C_{0}^{\infty}$ such that $0\leq\psi\leq 1$, $\psi$ has support where $|t|\leq 2$ and $\psi=1$ on [1,1][-1,1]. Let
(B.12) $G_{0}(t)=(2\pi)^{-1}\int\widehat{G}(\tau)\psi(\tau)\,d\tau$
(B.13) $G_{j}(t)=(2\pi)^{-1}\int\widehat{G}(\tau)\big(\psi(2^{-j}\tau)-\psi(2^{1-j}\tau)\big)\,d\tau\qquad j\in\mathbf{Z}_{+}$
Clearly, $G_{j}$ is analytic $N\times N$ matrix and that $G=\sum_{j=0}^{\infty}G_{j}$. We also have
(B.14) $|\tau^{\nu}\widehat{G}(\tau)|\leq\int|G(t)|+|G^{(\nu)}(t)|\,dt$
which gives
(B.15) $|G_{j}(t)|\leq C2^{j(1-\nu)}\int|G(s)|+|G^{(\nu)}(s)|\,ds\qquad\text{if $|\operatorname{Im}t|<\varepsilon_{j}=2^{-j}$}$
Thus we find from Lemma B.1 that
(B.16) $G_{j}(t)=Q^{{\varepsilon}_{j}}(t,B,G)P(t,B)+R^{{\varepsilon}_{j}}(B,G){\quad}\text{if}{\quad}\|B\|<1$
where
(B.17) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q(t,B,G)\|\leq C_{{\alpha}{\beta}}2^{-j(\nu-{\alpha}-|{\beta}|-3)}\int(\|G\|+\|G^{(\nu)}\|)\,dt$
$\|\partial_{B}^{{\beta}}R(B,G)\|\leq C_{{\beta}}2^{-j(\nu-|{\beta}|-2)}\int(\|G\|+\|G^{(\nu)}\|)\,dt$
With $\nu$ chosen as in (B.11) we find by summing up the convergent series, that
(B.18) $Q(t,B,G)=\sum_{j=0}^{\infty}Q^{{\varepsilon}_{j}}(t,B,G)\text{ and }R(t,B,G)=\sum_{j=0}^{\infty}R^{{\varepsilon}_{j}}(t,B,G)$
are smooth functions such that (B.4) and (B.11) hold. ∎
Now, we can prove the following symmetric version of the Malgrange division theorem.
Theorem B.3.
Let $G(t,x)\in C^{\infty}_{0}(\overline{\omega})$ with values in $N\times N$ matrices, where $\omega$ is a neighborhood of the origin in $\mathbf{R}\times\mathbf{R}^{n}$, then for any $M(x)\in C^{\infty}(\overline{\omega})$ with values in symmetric $N\times N$ matrices we find
(B.19) $G(t,x)=Q(t,x)P(t,M(x))+R(x)\qquad(t,x)\in\omega$
where QQ\/ and RR\/ are smooth $N\times N$ systems depending linearly on $G$ satisfying
(B.20) $\|\partial_{t}^{{\alpha}}\partial_{x}^{{\beta}}Q(t,x)\|\leq C_{{\alpha}{\beta}}\|G\|_{j}{\quad}j={\alpha}+|{\beta}|+4$
$\|\partial_{x}^{{\beta}}R(x)\|\leq C_{{\beta}}\|G\|_{k}{\quad}k=|{\beta}|+3$
Here $\|F\|_{j}=\sum_{|\alpha|+|\beta|\leq j}\|\partial_{t}^{{\alpha}}\partial_{x}^{{\beta}}F\|$ is the $L^{2}$ Sobolev norm.
Observe that Theorem B.3 gives that the mapping
$(Q(t,x),R(x))\mapsto G(t,x)=Q(t,x)P(t,M(x))+R(x)$
is a $C^{\infty}$ surjective tame mapping, having a $C^{\infty}$ tame right inverse of degree 4, see Appendix C, but that mapping is not tame for Malgranges preparation theorem in general. Observe that the symmetric scalar case when $N=1$ follows directly from the inverse function theorem.
Proof.
If $G(t,x)\in C^{\infty}_{0}(\mathbf{R}\times\mathbf{R}^{n})$, then by using Proposition A.2 we obtain (B.19) with $Q(t,x)=Q(t,M(x),G(\cdot,x))$ and $R(x)=R(M(x),G(\cdot,x))$. Since $Q(t,x)$ and $R(x)$ depend linearly on $G(t,x)$ we obtain (B.20) from (B.11) by Leibniz’ rule, since any $L^{1}$ Sobelev norm of test functions is bounded by the corresponding $L^{2}$ norm. ∎
In the preparation theorems, we have only done right symmetric division by matrix valued polynomials. By taking transposes we also obtain the results for left symmetric divisions, this does not affect the norms of the matrices.
Corollary B.4.
Let $G(t,x)\in C^{\infty}_{0}(\overline{\omega})$ with values in $N\times N$ matrices, where $\omega$ is a neighborhood of the origin in $\mathbf{R}\times\mathbf{R}^{n}$, then for any $M(x)\in C^{\infty}(\overline{\omega})$ with values in symmetric $N\times N$ matrices we can write
(B.21) $G(t,x)=P(t,M(x))Q(t,x)+R(x)\qquad(t,x)\in\omega$
where QQ\/ and RR\/ are smooth $N\times N$ systems depending linearly on $G$ satisfying (B.20).
We find that (B.21) gives a $C^{\infty}$ surjective tame map of degree 0 having a $C^{\infty}$ tame right inverse of degree 4, see Example C.13.
Appendix C Tame Fréchet spaces and the Nash–Moser theorem
In this section, we will recall the Nash–Moser Theorem and the theory of tame Fréchet spaces and tame maps from Sections II and III in [4]. Recall that a Fréchet space is a convex and complete vector space with topology defined by a countable number of seminorms, thus it is a complete metrizable vector space. But for the Nash–Moser Theorem one needs Fréchet spaces with a more refined structure.
Definition C.1.
A grading on a Fréchet space is a collection of seminorms $\{\|\cdot\|_{n},\ n\in\mathbf{N}\}$ that defines the topology and is increasing: $\|f\|_{j}\leq\|f\|_{k}$ when $j\leq k$.
Of course, every Fréchet space has a grading, for example by adding to each seminorm all the lower ones. Examples are $C^{\infty}$ functions and vector bundles with $C^{k}$ norms. Closed subspaces of graded spaces are graded with the induced norms, and Cartesian products $F\times G$ can be graded with the norms $\|(f,g)\|_{n}=\|f\|_{n}+\|g\|_{n}$.
Definition C.2.
Two gradings $\|\cdot\|_{n}$ and $\|\cdot\|^{\prime}_{n}$ are tamely equivalent of degree $r$ and base $b$ if
(C.1) $\|f\|_{n}\leq C_{n}\|f\|^{\prime}_{n+r}\quad\text{and}\quad\|f\|^{\prime}_{n}\leq C^{\prime}_{n}\|f\|_{n+r}$
for any $n\geq b$. If this holds for any $n$ then they are equivalent of degree $r$.
For example, if $X$ is a compact manifold, then the $C^{k}$ norms, Hölder norms and Sobolev norms are tamely equivalent.
Definition C.3.
A linear map $L:\ F\mapsto G$ of a graded space into another is tame, if it satisfies a tame estimate of degree $r$ and base $b$:
(C.2) $\|Lf\|_{n}\leq C_{n}\|f\|_{n+r}$
for any $n\geq b$. If (C.2) holds for any $n$ then $L$ is tame of degree $r$.
Example C.4.
A $C^{\infty}$ linear partial differential operator $L$ of order $m$ on $C^{\infty}$ is a tame of degree $m$ since $\|Lf\|_{n}\leq C_{n}\|f\|_{n+m}$ for the $C^{k}$ norms for any $n$. In particular, multiplication with $C^{\infty}$ functions are tame of degree 0. Solutions to invertible linear systems of smooth equations give tame maps of degree 0 by Leibniz’ rule. Similarly, for any $C^{\infty}$ system $M(x)$, we find that the mapping
$(Q(t,x),R(x))\mapsto G(t,x)=Q(t,x)(t\operatorname{Id}_{n}+M(x))+R(x)$
is tame of degree 0.
Example C.5.
The linear mapping $L:C^{\infty}[0,1]\mapsto C^{\infty}[-1,1]$ given by $Lf(x)=f(x^{2})$ is tame with $\|Lf\|_{n}\leq C_{n}\|f\|_{n}$. The image of $L$ is the closed subspace of even functions $C_{S}^{\infty}[-1,1]$. But the inverse $L^{-1}:\ C_{S}^{\infty}[-1,1]\mapsto C^{\infty}[0,1]$ is not tame, since the $2n$:th Taylor coefficient of $Lf=g$ is the $n$:th Taylor coefficient of $f=L^{-1}g$ so the best estimate is $\|L^{-1}f\|_{n}\leq C_{n}\|f\|_{2n}$.
Compositions of linear tame maps are linear and tame, and $L$ is a tame isomorphism if $L$ is linear isomorphism such that $L$ and $L^{-1}$ are tame.
Definition C.6.
Let $F$ and $G$ be graded spaces. Then we say that $F$ is a tame direct summand of $G$ if we can find tame linear maps $L:\ F\mapsto G$ and $M:\ G\mapsto F$ so that the composition ML:FFML:\ $\mapsto F$ is the identity: $F\overset{L}{\mapsto}G\overset{M}{\mapsto}F$.
Definition C.7.
We say that a graded space is tame if it is a tame direct summand of a space $\Sigma(B)$ of exponentially decreasing sequences in some Banach space $B$.
A Cartesian product of two tame spaces is tame.
Theorem C.8.
The space $C^{\infty}(X)$ is tame if $X$ is a compact manifold (with or without boundary), and if $V$ is a vector bundle over $X$ then the sections of the vector bundle $C^{\infty}(X,V)$ is also tame, since it can be written as a tame direct summand of a tame space.
Theorem C.9.
If $X$ is a compact manifold with boundary, then $C^{\infty}_{0}(X)$ and $C^{\infty}_{0}(X,V)$ are tame.
Next, we shall look at nonlinear maps.
Definition C.10.
Let $F$ and $G$ be graded spaces and $P:\ F\supseteq U\mapsto G$ a nonlinear map, then we say that $P$ satisfies a tame estimate of degree $r$ and base $b$ if
(C.3) $\|P(f)\|_{n}\leq C_{n}(1+\|f\|_{n+r})$
for all $f\in U$ and all $n\geq b$. If (C.3) holds for any $n$ then $P$ is tame of degree $r$. We say that $P$ is a tame map if $P$ is defined on an open set, is continuous and satisfies a tame estimate in a neighborhood of any point.
As before, if (C.3) holds for any $n$ then we say $P$ satisfies a tame estimate of degree $r$.
Theorem C.11.
A composition of tame maps is tame.
The following is the Nash–Moser theorem that we shall use.
Theorem C.12.
Let $F$ and $G$ be tame spaces and $P:\ F\supseteq U\mapsto G$ a $C^{\infty}$ tame map of the open set $U$. If the equation for the derivative $DP(f)h=k$ has a unique solution $h=VP(f)k$ for any $f\in U$ and any $k$, and the family of inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally invertible and each local inverse is a $C^{\infty}$ tame map.
If the derivative $DP(f)h=k$ has a solution $h=VP(f)k$ for any $f\in U$ and any $k$, and the family of right inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally surjective and each local right inverse is a $C^{\infty}$ tame map.
If the derivative $DP(f)$ is injective for any $f\in U$ and the family of left inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally injective and each local left inverse is a $C^{\infty}$ tame map.
Example C.13.
It follows from Corollary B.4 that the mapping
$C^{\infty}(\mathbf{R}\times\mathbf{R}^{n})\times C^{\infty}(\mathbf{R}^{n})\ni(Q(t,x),M(x))\mapsto G(t,x)=P(t,M(x))Q(t,x)+R(x)$
with $M(x)=M^{*}(x)\ \forall\,x$ is a $C^{\infty}$ tame surjective map of degree 0 with local $C^{\infty}$ tame right inverses of degree 4.
The proofs of these results and more examples and counterexamples can be found in sections II and III in [4].
Acknowledgment.
I would like to thank Gregory Berkolaiko at Texas A&M, who asked me about symmetric preparation theorems like (1.4) which started this project. We then worked together on the formal power series solution in Section 3. I am also indebted to Dmitry Kerner at Ben-Gurion University, who informed us about the nested Artin Approximation, which gives analytic solutions in the two dimensional case from the formal power series solution, see Remark 2.2.

### v3

@@
+arXiv:2601.12458v3[math.CA] 28 Feb 2026
+arXiv:2601.12458v3[math.CA] 28 Feb 2026
 Symmetric Preparation of Systems
-Abstract.
-In this paper we generalize the Weierstrass and Malgrange preparation theorems to the symmetric matrix valued case, proving symmetric preparation of analytic and smooth symmetric systems that vanishes of first order.
-2010 Mathematics Subject Classification:
-Primary 32B05 Secondary 47A56, 26B40
+Nils Dencker
+Lund University
+Department of Mathematics, University of Lund, Box 118, S-221 00 Lund, Sweden nils.dencker@gmail.com
 1. Introduction and statement of results
-If $f$ is $C^{\infty}$ satisfying $\partial_{t}f(0,0)\neq 0$ and $f(0,0)=0$ then by the Malgrange preparation theorem there exist $C^{\infty}$ functions $c(t,x)$ and $r(x)$ such that $c(0,0)\neq 0$, $r(0)=0$ and
-(1.1) $f(t,x)=c(t,x)\left(t+r(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
+If $f$ is $C^{\infty}$ satisfying $\partial_{t}^{m}f(0,0)\neq 0$ and $\partial_{t}^{j}f(0,0)=0$ for $j<m$, then by the Malgrange preparation theorem there exist $C^{\infty}$ functions $c(t,x)$ and $r(t,x)=\sum_{j=0}^{m-1}t^{j}r_{j}(x)$ such that $c(0,0)\neq 0$, $r_{j}(0)=0,\ \forall\,j$ and
+(1.1) $f(t,x)=c(t,x)\left(t^{m}+r(t,x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
 locally near $(0,0)$.
-In the analytic case, the Weierstrass preparation theorem, one has uniqueness but not in the $C^{\infty}$ case, except for real valued $f(x)$ when one can use the implicit function theorem.
-This also holds for higher order vanishing of $f$ and there are corresponding division results.
-The Malgrange preparation theorem has been generalized to $N\times N$ systems of functions in [3].
-The simplest example is for smooth $N\times N$ system $F(t,x)$ satisfying $|\partial_{t}F(0,0)|\neq 0$ and $F(0,0)=0$.
-Then there exist smooth $N\times N$ systems $C(t,x)$ and $M(x)$ such that $|C(0,0)|\neq 0$, $M(0)=0$ and
-(1.2) $F(t,x)=C(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
-locally near $(0,0)$, see [2]. In the analytic case, one has uniqueness but not in the smooth case.
+In the analytic case, the Weierstrass preparation theorem, one has uniqueness but not in the $C^{\infty}$ case, except for $m=1$ and real valued $r(t,x)$ when one can use the implicit function theorem.
+There are also corresponding division results.
+The Malgrange preparation theorem has been generalized to $N\times N$ systems of functions in [3]. An example is for smooth $N\times N$ system $F(t,x)$ satisfying
+(1.2) $|\partial_{t}^{m}F(0,0)|>0\quad\text{and}\quad\partial_{t}^{j}F(0,0)=0,\quad j<m$
+Then there exist smooth $N\times N$ systems $C(t,x)$ and $M_{j}(x)$, $j<m$, such that $|C(0,0)|\neq 0$, $M_{j}(0)=0$, $\forall\,j$, and
+(1.3) $F(t,x)=C(t,x)\left(t^{m}\operatorname{Id}_{N}+\sum_{j=0}^{m-1}t^{j}M_{j}(x)\right)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
+locally near $(0,0)$, this is the special case of [3, Th. 5.3]. In the analytic case, one has uniqueness but not in the smooth case.
 We also proved the following division theorem in [3].
-If $F(t,x)$ is a smooth $N\times N$ system, then for any smooth $N\times N$ system $M(x)$ there exist smooth $N\times N$ systems $C(t,x)$ and $R(x)$ such that
-(1.3) $F(t,x)=C(t,x)\left(t\operatorname{Id}+M(x)\right)+R(x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
-locally near $(0,0)$.
+If $F(t,x)$ is a smooth $N\times N$ system satisfying (1.2), then for any smooth $N\times N$ system $G(t,x)$ there exist smooth $N\times N$ systems $C(t,x)$ and $R_{j}(x)$, $j<m$, such that
+(1.4) $G(t,x)=C(t,x)F(t,x)+\sum_{j=0}^{m-1}t^{j}R_{j}(x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
+locally near $(0,0)$, this is a special case of [3, Th.
+5.9].
 As before, one has uniqueness in the analytic case but not in the smooth case.
-We also have preparation results for higher order vanishing of $F$, adjoint preparation results and corresponding division results, see [3].
+By taking the adjoints, we also obtain adjoint preparation results and division results, see [3].
 Now, the question about symmetric preparation results was raised by Gregory Berkolaiko [1].
 The simplest case is for smooth (or analytic) symmetric $N\times N$ matrices $F(t,x)$ satisfying $\partial_{t}F(0,0)>0$ and $F(0,0)=0$.
-Then, does it exist smooth (or analytic) $N\times N$ matrices $U(t,x)$ and $M(x)$, such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric with $M(0)=0$ satisfying
+Then the question was: does it exist smooth (or analytic) $N\times N$ matrices $U(t,x)$ and $M(x)$, such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric with $M(0)=0$, satisfying
 (1.5) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)\qquad(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$
 near $(0,0)$, and is this preparation unique?
-Observe that the conditions on $F(t,x)$ are necessary for (1.4) to hold since $U(0,0)U^{*}(0,0)=\partial_{t}F(0,0)>0$ and $F(0,0)=U(0,0)M(0)U(0,0)^{*}=0$.
-Here and in the following we say that the matrix $A$ is symmetric if $A^{*}=A$ and antisymmetric if $A^{*}=-A$, where $A^{*}=\overline{A}^{t}$.
+Observe that the conditions on $F(t,x)$ are necessary for (1.5) to hold since $U(0,0)U^{*}(0,0)=\partial_{t}F(0,0)>0$ and $F(0,0)=U(0,0)M(0)U^{*}(0,0)$.
+Here and in the following we say that the matrix $A$ is symmetric if $A^{*}=A$ and antisymmetric if $A^{*}=-A$, where $A^{*}=\overline{A}^{t}$, and when we say that $A>0$ then we assume that $A$ is symmetric.
+This result was proved in [4], and in this paper we shall generalize to the higher order case.
+The following is the main result of the paper.
+Theorem 1.1.
+Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ satisfying
+(1.6) $\partial_{t}^{m}F(0,0)>0\quad\text{and}\quad\partial_{t}^{j}F(0,0)=0,\quad j<m$
+then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M_{j}(x)$, $j<m$, such that $|U(0,0)|\neq 0$, $M_{j}(x)$ is symmetric such that $M_{j}(0)=0,\ \forall\,j,$ and
+(1.7) $F(t,x)=U(t,x)\left(t^{m}\operatorname{Id}_{N}+\sum_{j=0}^{m-1}t^{j}M_{j}(x)\right)U^{*}(t,x)\quad\text{in $\omega$}$
+If $F(t,x)$ is analytic then we obtain (1.7) with analytic $U(t,x)$ and ${\mathbb{M}}(t,x)$. One can choose $U>0$ and then the preparation is unique.
 Remark 1.2.
 Observe that the preparation (1.7) is not unique in general.
-In fact, if $U(t,x)$ and M(x) are solutions to (1.4), then for any orthogonal $A(x)=(A^{-1})^{*}(x)$ we find that $\widetilde{U}(t,x)=U(t,x)A(x)$ and $\widetilde{M}(x)=A^{*}(x)M(x)A(x)$ solves (1.4).
-Actually, the preparation is unique in the case when $U$ is symmetric and $U>0$, see Theorem 1.2. We are going to prove the following, which is the main result of the paper.
-Theorem 1.2.
-Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ such that
-(1.5) $\partial_{t}F(0,0)>0\quad\text{and}\quad F(0,0)=0$
-Then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M(x)$ such that $|U|\neq 0$, $M$ is symmetric , $M(0)=0$ and
-(1.6) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)\qquad\text{in $\omega$}$
-If $F(t,x)$ is analytic then we obtain (1.6) with analytic $U(t,x)$ and $M(x)$. This preparation is unique if $U$ is symmetric and $U>0$.
-Remark 1.3.
-Observe that we also get uniqueness in other cases, for example if $U<0$ by taking U-U for for the unique preparations with $U>0$.
-The proof of Theorem 1.2 also gives the preparation (1.6) for some $0\neq|F(0,0)|\ll 1$ with $|M(0)|\ll 1$, but the bounds depend on the (semi)norms of $F$.
-If $F(0,0)\neq 0$ then by subtracting $F(0,0)$ we obtain (1.6) with the remainder $F(0,0)$ giving the following result.
+For example, if $U(t,x)$ and ${\mathbb{M}}(t,x)$ are solutions to (1.7), then for any orthogonal $A(x)=(A^{-1})^{*}(x)$ we find that $\widetilde{U}(t,x)=U(t,x)A(x)$ and $\widetilde{\mathbb{M}}(t,x)=A^{*}(x){\mathbb{M}}(t,x)A(x)$ solves (1.7).
+But if $U(t,x)>0$ and $UA(t,x)>0$ then $A(t,x)\equiv\operatorname{Id}_{N}$.
+In fact, if $U>0$ then $U$ has a ON base of eigenvectors, and if $UA$ is symmetric then $A^{*}U=UA$ which gives $U=AUA$.
+Then $A$ preserves the eigenspace with the largest eigenvalue $\|U\|$ and $A^{2}=\operatorname{Id}$ on this space.
+By iteration over decreasing eigenvalues of $U$ we find $A^{2}=\operatorname{Id}_{N}$.
+Then $A=A^{-1}=A^{*}$ has eigenvalues $\pm 1$, and if $UA>0$ there cannot be any negative eigenvalues, thus $A=\operatorname{Id}_{N}$.
+Now (1.7) gives that $\partial_{t}^{j}F(0,0)=0$ for $j<m$ and $m!U(0,0)U^{*}(0,0)=\partial_{t}^{m}F(0,0)>0$, so the conditions are necessary.
+If $\partial_{t}^{j}F(0,0)\neq 0$ for $j<m$ then by subtracting suitable terms we obtain the following result.
 Corollary 1.3.
-Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ such that
-(1.7) $\partial_{t}F(0,0)>0$
-Then there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and $M(x)$ such that $|U(0,0)|\neq 0$, $M(x)$ is symmetric $\forall\,x$, $M(0)=0$ and
-(1.8) $F(t,x)=U(t,x)\left(t\operatorname{Id}_{N}+M(x)\right)U^{*}(t,x)+F(0,0)\qquad\text{in $\omega$}$
-If $F(t,x)$ is analytic then we obtain (1.8) with analytic $U(t,x)$ and $M(x)$. This preparation is unique if $U>0$.
-Note that the uniqueness only holds for symmetric preparations, since Remark 1.1 gives counterexamples for nonsymmetric $U$.
+Assume that $F(t,x)$ is a smooth symmetric $N\times N$ matrix valued function of $(t,x)\in\mathbf{R}\times\mathbf{R}^{n}$ satisfying
+(1.8) $\partial_{t}^{m}F(0,0)>0$
+Then for any smooth symmetric $N\times N$ system $\mathbb{R}(t,x)$ such that $\partial_{t}^{j}\mathbb{R}(0,0)=\partial_{t}^{j}F(0,0)$, $j<m$, and $\partial_{t}^{m}\mathbb{R}(0,0)<\partial_{t}^{j}F(0,0)$, there exists a neighborhood $\omega$ of $(0,0)$ and smooth $N\times N$ matrix valued functions $U(t,x)$ and ${\mathbb{M}}(t,x)$ such that $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, and
+(1.9) $F(t,x)=U(t,x)\left(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x)\right)U^{*}(t,x)+\mathbb{R}(t,x)\qquad\text{in $\omega$}$
+If $F(t,x)$ is analytic then we obtain (1.9) with analytic $U(t,x)$, ${\mathbb{M}}(t,x)$ and $\mathbb{R}(t,x)$. One can choose $U>0$ and then the preparation is unique.
+Note that the uniqueness does not hold in general, since Remark 1.2 gives counterexamples for nonsymmetric $U$.
 Since (1.7) is not linear in $U(t,x)$ the proof approaches of Weierstrass or Malgrange preparation results do not work in the symmetric case, instead we shall use inverse function theorems.
@@
 2. Formal power series solutions
-If $F$ is analytic we may solve (1.4) by taking the Taylor expansion of $F(t,x)$:
-(2.1) $F(t,x)=\sum_{{j,\alpha}}F_{j,\alpha}t^{j}x^{\alpha}$
-where $F_{0,0}=F(0,0)=0$ and $F_{1,0}=\partial_{t}F(0,0)>0$, and take the formal power series expansions $U(t,x)=\sum_{{j,\alpha}}U_{\begin{subarray}{c}j,\alpha\end{subarray}}t^{j}x^{\alpha}$ where $|U_{0,0}|\neq 0$ and $M(x)=\sum_{\alpha\neq 0}M_{\alpha}x^{\alpha}$ is symmetric such that $M_{0}=M(0)=0$.
-Then (1.4) gives that
-(2.2) $\sum_{{j,\alpha}}F_{j,\alpha}t^{j}x^{\alpha}=\sum_{{j,\alpha}}U_{j,\alpha}t^{j}x^{\alpha}\left(t\operatorname{Id}_{N}+\sum_{\gamma\neq 0}M_{\gamma}x^{\gamma}\right)\sum_{{k,\beta}}U_{k,\beta}^{*}t^{k}x^{\beta}$
-Simplifying the notations so that $\sum_{{j,\alpha}}A_{\begin{subarray}{c}j,\alpha\end{subarray}}t^{j}x^{\alpha}=\sum_{j,k}A_{j,k}t^{j}x^{k}$, we get
-(2.3) $\sum_{j,k}F_{j,k}t^{j}x^{k}=\sum_{j,k}U_{j,k}t^{j}x^{k}\left(t\operatorname{Id}_{N}+\sum_{m>0}M_{m}x^{m}\right)\sum_{{i,\ell}}U_{i,\ell}^{*}t^{i}x^{\ell}$
+We shall assume that $F(t,x)$ is given by a formal power series expansion:
+(2.1) $F(t,x)=\sum_{{j,k}}F_{j,\alpha}t^{j}x^{k}$
+where $F_{j,0}=\partial_{t}^{j}F(0,0)/j!=0$ for $j<m$ and $F_{m,0}=\partial_{t}^{m}F(0,0)/m!>0$.
+We shall take the formal power series expansions $U(t,x)=\sum_{{j,k}}U_{\begin{subarray}{c}j,k\end{subarray}}t^{j}x^{k}$ where $|U_{0,0}|\neq 0$ and
+(2.2) ${\mathbb{M}}(t,x)=\sum_{j<m}t^{j}{\mathbb{M}}_{j}(x)=\sum_{\begin{subarray}{c}j<m\\ k>0\end{subarray}}M_{j,n}t^{j}x^{k}$
+is symmetric such that ${\mathbb{M}}_{j}(0)=M_{j,0}=0$, $j<m$.
+Then (1.7) gives
+(2.3) $\sum_{j,k}F_{j,k}t^{j}x^{k}=\sum_{j,k}U_{j,k}t^{j}x^{k}\Big(t^{m}\operatorname{Id}_{N}+\sum_{\begin{subarray}{c}j,k\end{subarray}}M_{j,k}t^{j}x^{k}\Big)\sum_{{j,k}}U_{j,k}^{*}t^{j}x^{k}$
 and by matching coefficients we obtain that
-(2.4) $F_{j,k}=\sum_{\begin{subarray}{c}i<j\\ \ell\leq k\end{subarray}}U_{i,\ell}U_{j-1-i,k-\ell}^{*}+\sum_{\begin{subarray}{c}i\leq j\\ \ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{i,\ell}M_{m}U_{j-i,k-m-\ell}^{*}$
-The first terms of this expansion are $F_{0,0}=0$, $F_{0,1}=U_{0,0}M_{1}U_{0,0}^{*}$ and $F_{1,0}=U_{0,0}U_{0,0}^{*}$ which gives that $|U_{0,0}|\neq 0$ and $M_{1}=U_{0,0}^{-1}F_{0,1}(U_{0,0}^{-1})^{*}$.
-When $j=0$ we obtain from (2.4) that
-(2.5) $F_{0,k}=\sum_{\begin{subarray}{c}\ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}\qquad\forall\,k$
-so recursion gives
-(2.6) $U_{0,0}M_{k}U_{0,0}^{*}=F_{0,k}-\sum_{\begin{subarray}{c}0\leq\ell\leq k-m\\ 0<m<k\end{subarray}}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}\qquad\forall\,k$
-We then obtain symmetric $M_{j}$ by
-(2.7) $M_{k}=U_{0,0}^{-1}F_{0,k}(U_{0,0}^{-1})^{*}-\sum_{\begin{subarray}{c}0\leq\ell\leq k-m\\ 0<m<k\end{subarray}}U_{0,0}^{-1}U_{0,\ell}M_{m}U_{0,k-m-\ell}^{*}(U_{0,0}^{-1})^{*}\qquad\forall\,k$
-so $M_{k}$ is determined by $F_{0,k}$, $U_{0,\ell}$ with $\ell<k$ and $M_{m}$ with $m<k$.
-In general, we obtain from (2.4) with $j$ replaced by $j+1$ that
-(2.8) $2\operatorname{Re}U_{j,k}U_{0,0}^{*}=F_{j+1,k}-\sum_{\begin{subarray}{c}0<i<j\\ \ell\leq k\end{subarray}}U_{i,\ell}U_{j-1-i,k-\ell}^{*}-\sum_{\begin{subarray}{c}i\leq j+1\\ \ell\leq k-m\\ 0<m\leq k\end{subarray}}U_{i,\ell}M_{m}U_{j+1-i,k-m-\ell}^{*}$
-where the right hand side is determined by $F_{j+1,k}$, $M_{m}$ with $m\leq k$ determined by (2.6) and $U_{i,\ell}$ with $i<j$ and $\ell\leq k$ or $i\leq j+1$ and $\ell\leq k-1$, which gives $i+\ell<j+k$ and $\ell\leq k$ or $i+\ell\leq j+k$ and $\ell\leq k-1$.
-Since $|U_{0,0}|\neq 0$, (2.8) can be solved to give $U_{j,k}$ and by Lemma A.1 we get a unique symmetric solution $U_{j,k}$ when $U_{0,0}$ is symmetric. For $k=0$ we find from (2.8) that
-(2.9) $2\operatorname{Re}U_{j,0}U_{0,0}^{*}=F_{j+1,0}-\sum_{\begin{subarray}{c}0<i<j\end{subarray}}U_{i,0}U_{j-1-i,0}^{*}$
-which gives $2\operatorname{Re}U_{1,0}U_{0,0}^{*}=F_{2,0}$.
-Since $|U_{0,0}|\neq 0$ we obtain by induction $U_{j,0}$ for any $j$ from (2.9), and if $U_{0,0}$ is symmetric then Lemma A.1 gives a unique symmetric $U_{j,0}$.
-Thus, one can determine $U_{j,k}$ and $M_{k}$ from (2.7) and (2.8) by induction over $j+k=p\geq 0$.
-For $p=0$ we can take any $U_{0,0}$ such that $U_{0,0}U_{0,0}^{*}=F_{1,0}>0$.
-Then for fixed $p>0$ we get the values of $U_{p,0}$ from (2.9) so we can do a recursion over (pr,r)(p-r,r) for $r=0$ to $p$.
-If we have determined $U_{j,k}$ for $j+k<p$ and $U_{p-\ell,\ell}$ for $\ell\leq r$ for some $r<p$ then by (2.8) we find that $U_{p-r-1,r+1}$ only depends on $F_{p-r,r+1}$, $U_{i,\ell}$ for $i+\ell\leq p$ and $\ell\leq r$ or $i+\ell<p$ and $\ell\leq r+1$, which are previously determined, and on $M_{m}$ with $m\leq r+1$ which by (2.7) are given by $F_{0,m}$ and $U_{0,\ell}$ with $\ell\leq r<p$ which also have been determined.
-By recursion over $r$ and induction over $p$ we get the expansion, and if $U_{0,0}$ is symmetric we get unique symmetric $U_{j,k}$ by (2.8) and Lemma A.1, which gives the following result.
+(2.4) $F_{j,k}=\sum_{i,\ell}U_{i,\ell}U_{j-m-i,k-\ell}^{*}+\sum_{\begin{subarray}{c}i,\ell,p,q\end{subarray}}U_{i,\ell}M_{p,q}U_{j-p-i,k-q-\ell}^{*}$
+where we sum over the nonnegative index.
+The first terms of this expansion are $F_{0,0}=U_{0,0}M_{0,0}U_{0,0}^{*}=0$, $F_{0,1}=U_{0,0}M_{0,1}U_{0,0}^{*}$, $F_{j,0}=0$, $j<m$, and $F_{m,0}=U_{0,0}U_{0,0}^{*}$.
+Here we can choose $U_{0,0}=(F_{m,0})^{1/2}$ which is the unique positive definite symmetric solution.
+In fact, since $U_{0,0}$ is invertible we can write any other positive symmetric solution as $U_{0,0}A$ and then $(U_{0,0}A)^{*}=A^{*}U_{0,0}$.
+Since $(U_{0,0}A)^{2}=U_{0,0}^{2}$, $A$ has to be orthogonal.
+But then Remark 1.2 gives that $A=\operatorname{Id}_{N}$.
+Thus, we get a unique $U_{0,0}>0$.
+When $j<m$ we obtain from (2.4) that
+(2.5) $U_{0,0}M_{j,k}U_{0,0}^{*}=F_{j,k}-\sum_{\begin{subarray}{c}i+\ell<j+k\end{subarray}}U_{i,\ell}M_{p,q}U_{j-p-i,k-q-\ell}^{*}$
+since $|U_{0,0}|\neq 0$ we obtain symmetric $M_{j,k}$ that is given by $F_{j,k}$, $U_{i,\ell}$ with $i\leq j$ and $\ell\leq k$, and by $M_{p,q}$ with $p\leq j$, $q\leq k$ and $p+q<j+k$, Thus $M_{j,k}$ is determined by $F_{j,k}$, $U_{i,\ell}$ with $\ell<k$ or $\ell=k$ and $i<j$ and by $M_{p,q}$ with $q<k$ or $q=k$ and $p<j$.
+With $j$ replaced by $j+m$, we also obtain from (2.4) that
+(2.6) $2\operatorname{Re}U_{j,k}U_{0,0}^{*}=F_{j+m,k}-\sum_{\begin{subarray}{c}0<i+\ell<j+k\end{subarray}}U_{i,\ell}U_{j-i,k-\ell}^{*}-\sum_{q>0}U_{i,\ell}M_{p,q}U_{j+m-p-i,k-q-\ell}^{*}$
+where the right hand side is given by $F_{j+m,k}$, $U_{i,\ell}$ with $i\leq j+m$ and $\ell<k$ or with $i+\ell<j+k$, thus by $\ell<k$ or $\ell=k$ and $i<j$, and $M_{p,q}$ with $p<m$ and $q\leq k$.
+Since $U_{0,0}$ is invertible, we can solve (2.6).
+In the case $U_{0,0}>0$, Lemma A.1 gives a unique symmetric $U_{j,k}$.
+We can now determine $U_{j,k}$ and $M_{j,k}$ from (2.5) and (2.6) by induction over $j\geq 0$ for fixed $k\geq 0$ starting with $U_{0,k}$ and $M_{0,k}$ which we first determine by induction over $k$.
+Since $M_{0,0}=0$ we start with $U_{0,0}$ and $M_{0,1}=U_{0,0}^{-1}F_{0,1}(U_{0,0}^{*})^{-1}$, and assume that $F_{j,k}$ is known.
+We assume that $U_{0,\ell}$ and $M_{0,\ell+1}$ is known for $\ell<k$.
+Then $U_{0,k}$ is determined by these and then $M_{0,k+1}$ is determined by $U_{0,\ell}$ and $M_{0,\ell}$ for $\ell\leq k$.
+By induction we obtain $U_{0,k}$ and $M_{0,k+1}$ for any $k\geq 0$.
+Then we use induction over $j\geq 0$ for $U_{j,k}$ and $M_{j,k+1}$ starting with $k=0$.
+We assume that we have determined $U_{i,\ell}$ and $M_{i,\ell+1}$ when $\ell<k$ and when $\ell=k$ for $i<j$.
+Then $U_{j,k}$ is determined by these and we then obtain $M_{j,k+1}$ from $U_{i,\ell}$ when $\ell\leq k$ and from $M_{i,\ell+1}$ when $\ell<k$ and when $\ell=k$ for $i<j$.
+By induction we obtain $U_{j,k}$ and $M_{j,k+1}$ for any $j$ and $k$.
+Observe that any finite expansion of $U(t,x)$ is positive definite if $|t|+|x|\ll 1$.
 Theorem 2.1.
-If $F(t,x)$ is analytic and symmetric such that $F(0,0)>0$ and $F(0,0)=0$ then there exists formal power series $U(t,x)$ and $M(x)$ solving (1.4) modulo $O(|t|^{N}+|x|^{N})$, $\forall\,N$.
-We get a unique power series solution in the case when $U$ is symmetric.
-But is it not clear that the formal power expansion converge to an analytic function.
-Remark 2.2.
-If $n=1$ and $F$ is analytic, then by  [1] the nested Artin approximation (see e.g.
-Theorem 5.18 in [6]) gives analytic $\overline{U}(x,y)=\sum_{j,k}\overline{U}_{jk}t^{j}x^{k}$ and $\overline{M}(x)=\sum_{k}\overline{M}_{k}x^{k}=\overline{M}^{*}(x)$ that solves
-$F(t,x)=\overline{U}(t,x)(tId+\overline{M}(x))\overline{U}^{*}(t,x)\qquad(t,x)\in\mathbf{R}^{2}$
-which gives Theorem 1.2 in the two dimensional analytic case.
-One also obtains that any finite number of terms in the expansions of $\overline{U}(t,x)$ and $\overline{M}(x)$ coincide with the corresponding terms in formal power series solutions $U(t,x)$ and $M(x)$.
-When $n>1$ this does not work in general, see the counterexamples in [6]. Instead we will use the inverse function theorem, see Section 3.
+If $F(t,x)$ is analytic (or formal power series) and symmetric such that $\partial_{t}^{m}F(0,0)>0$ and $\partial_{t}^{j}F(0,0)=0$, $\forall\,j<m$, then there exists formal power series $U(t,x)$ and ${\mathbb{M}}(t,x)$ given by (2.2) solving (1.7) modulo $O(|t|^{N}+|x|^{N})$, $\forall\,N$.
+We get a unique symmetric power series solution in the case when $U_{0,0}>0$.
+But is it not clear that in the analytic case, the formal power expansion converges to an analytic function. Instead we will use the inverse function theorem, see Section 3.
 3. Proof of Theorem 1.1 in the analytic case
-It is clear that $N\times N$ analytic systems form a Banach space with the usual $L^{{}^{\infty}}$ norm, and they can be localized to neighborhoods.
-We shall prove that if $U(t,x)$ and $M(x)$ are analytic such that $|U|\neq 0$, $M$ is symmetric and $M(0)=0$, then the mapping
-(3.1) $(U(t,x),M(x))\overset{\mathcal{F}}{\longrightarrow}\Big(\partial_{t}\big(U(t,x)(t\operatorname{Id}_{N}+M(x))U^{*}(t,x)\big),U_{0}(x)M(x)U_{0}^{*}(x)\Big)\\ =(F_{1}(t,x),F_{0}(x))$
-with $U_{0}(x)=U({0,x})$, is analytic and surjective in some neighborhood of $(0,0)$ for analytic and symmetric $F_{j}$, j=0, 1j=0,\,1, when $|F_{1}(0,0)|\neq 0$ and $F_{0}(0)=0$.
-If we obtain a solution to (3.1) in a neighborhood of $(0,0)$ then we obtain (1.6) with
-(3.2) $F(t,x)=F_{0}(x)+\int_{0}^{t}F_{1}(s,x)\,ds\qquad\text{near $(0,0)$}$
-thus $F_{0}(x)=F(0,x)$ and $F_{1}(t,x)=\partial_{t}F(t,x)$.
-When $M\equiv 0$ we find that $\partial_{t}(tUU^{*})=F_{1}$ so that
-(3.3) $U(t,x)U^{*}(t,x)=\frac{1}{t}\int_{0}^{t}F_{1}(s,x)\,ds\to F_{1}(0,x)\text{ when }t\to 0$
-For any such solution $U(t,x)$, we find that the differential of $\mathcal{F}$ at $(U,0)$ is given by
-(3.4) $(u(t,x),m(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}\big(2t\operatorname{Re}U(t,x)u^{*}(t,x)+U(t,x)m(x)U^{*}(t,x)\big),U_{0}(x)m(x)U_{0}^{*}(x)\Big)\\ =(A_{1}(t,x),A_{0}(x))$
-where $u$ and $m$ are analytic and $m=m^{*}$, which gives analytic symmetric $A_{j}$, j=0, 1j=0,\,1.
-We shall show that this differential is surjective in a neighborhood of $(0,0)$.
-We find from (3.4) that $m=U_{0}^{-1}A_{0}(U_{0}^{-1})^{*}$ and
-$2t\operatorname{Re}U(t,x)u^{*}(t,x)=\int_{0}^{t}A_{1}(s,x)\,ds+R(x)-U(t,x)m(x)U^{*}(t,x)$
-for some analytic symmetric $R(x)$. Since the right hand side has to vanish at $t=0$, we take $R(x)=U_{0}(x)m(x)U_{0}^{*}(x)=A_{0}(x)$ and then
-$2\operatorname{Re}U(t,x)u^{*}(t,x)=\frac{1}{t}\left(\int_{0}^{t}A_{1}(s,x)\,ds+A_{0}(x)-U(t,x)m(x)U^{*}(t,x)\right)=A(t,x)$
+It is clear that $N\times N$ analytic systems form a Banach space with the usual $L^{\infty}$ norm, and they can be localized to neighborhoods. We shall use the Taylor expansion notation
+$T_{k}(F)(t,x)=\sum_{j=0}^{k-1}t^{j}\partial_{j}^{j}F(0,x)/j!$
+and also the $k$ times integrations $I_{k}(F)(t,x)$ in $t$ such that $I_{1}(F)(t,x)=\int_{0}^{t}F(s,x)\,ds$ and $I_{k+1}(F)(t,x)=I_{1}(I_{k}(F))(t,x)$.
+Thus we find $\partial_{t}^{j}I_{k}(F)(0,x)\equiv 0$ when $j<k$ and $\partial_{t}^{k}I_{k}(F)(t,x)\equiv F(t,x)$.
+Observe that $t^{-k}I_{k}(F)(t,x)$ is analytic or smooth if $F(t,x)$ is.
+We shall prove that if $U(t,x)$ and ${\mathbb{M}}(t,x)$ are analytic such that $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)/j!$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, then the mapping
+(3.1) $(U(t,x),{\mathbb{M}}(t,x))\overset{\mathcal{F}}{\longrightarrow}\\ \Big(\operatorname{Re}\partial_{t}^{m}\big(U(t,x)(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x))U^{*}(t,x)\big),T_{m}(U{\mathbb{M}}U^{*})(t,x))\Big)\\ =(F_{1}(t,x),F_{0}(t,x))$
+is analytic and surjective in some neighborhood of $(0,0)$ for analytic and symmetric $F_{j}$, j=0, 1j=0,\,1, satisfying $|F_{1}(0,0)|\neq 0$, $\partial_{t}^{m}F_{0}\equiv 0$ and $F_{0}(t,0)\equiv 0$.
+If $\mathcal{F}(U,{\mathbb{M}})=(F_{1},F_{0})$ by (3.1) in a neighborhood of $(0,0)$ then we obtain (1.7) with
+(3.2) $F(t,x)=F_{0}(t,x)+I_{m}(F_{1})(t,x)\qquad\text{near $(0,0)$}$
+since $F_{0}(t,x)=T_{m}F(t,x)$ and $F_{1}(t,x)=\partial_{t}^{m}F(t,x)$.
+When ${\mathbb{M}}\equiv 0$ we find that $F_{0}\equiv 0$ and $\partial_{t}^{m}(t^{m}UU^{*})=m!UU^{*}=F_{1}$ when $t=0$ so $U(t,x)$ must be invertible close to $(0,0)$.
+For any such $U(t,x)$, we find that the differential of $\mathcal{F}$ at $(U,0)$ is given by
+(3.3) $({\bf u}(t,x),{{\bf m}}(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(2t^{m}\operatorname{Re}U(t,x){\bf u}^{*}(t,x)+U(t,x){{\bf m}}(t,x)U^{*}(t,x)\big),T_{m}(U{\bf m}U^{*})(t,x)\Big)\\ =(A_{1}(t,x),A_{0}(t,x))$
+where ${\bf u}(t,x)$ and ${\bf m}(t,x)$ are analytic such that $\bf m$} is symmetric, $\partial_{t}^{m}{\bf m}\equiv 0$ and ${\bf m}(t,0)\equiv 0$, which gives analytic symmetric $A_{j}$, j=0, 1j=0,\,1, such that $\partial_{t}^{m}A_{0}\equiv 0$ and $A_{0}(t,0)\equiv 0$.
+We shall show that this differential is surjective for such systems in a neighborhood of $(0,0)$. We find from (3.3) that ${\bf m}=T_{m}(U^{-1}A_{0}(U^{-1})^{*})$ and
+$2t^{m}\operatorname{Re}U(t,x){\bf u}^{*}(t,x)=I_{m}(A_{1})(t,x)+\mathbb{R}(t,x)-U{\bf m}U^{*}(t,x)$
+for some analytic symmetric $\mathbb{R}(t,x)$ such that $\partial_{t}^{m}\mathbb{R}\equiv 0$.
+Since the right hand side has to vanish of order $m$ at $t=0$, we find that $\mathbb{R}(t,x)=T_{m}(U{\bf m}U^{*})(t,x)=A_{0}(t,x)$ and then
+$2\operatorname{Re}U(t,x){\bf u}^{*}(t,x)=t^{-m}\Big(I_{m}(A_{1})(t,x)+A_{0}(t,x)-U{\bf m}U^{*}(t,x)\Big)=A(t,x)$
 which is analytic and symmetric.
-By substituting $m$ in $A$ we find that $A$ depends linearly on $A_{j}$, j=0, 1j=0,\,1, and there exists an analytic and antisymmetric $B(t,x)$ so that $2Uu^{*}=A+B$, which gives the solution
-(3.5) $u(t,x)=\frac{1}{2}(A(t,x)-B(t,x))(U^{-1})^{*}(t,x)$
-Thus (3.4) is surjective so by the inverse function theorem we obtain a local right inverse to $\mathcal{F}$ and local surjectivity of the nonlinear mapping (3.1) in a neighborhood $V$ of $(U,0)$.
-This neighborhood depend on $U(t,x)$ and thus by $F(t,x)$ by (3.3).
-Observe that $\mathcal{F}$ will also be surjective on some $(F_{1},F_{0})$ such that $F_{0}(0)\neq 0$.
-For any symmetric $M(x)$ such that $M(0)=0$, there exists a neighborhood $\omega$ of $(0,0)$ such that $(U,M)\in V$ at $\omega$, and then (3.5) gives a solution to (3.1) in $\omega$.
-But (3.4) is not injective, since different antisymmetric $B$ gives different solutions by (3.5).
-Actually, we have for fixed $U$ and $M\equiv 0$ that the mapping $(A_{0},A_{1},B)\mapsto(u,m)$ is bijective near $(0,0)$.
-It remains to prove uniqueness in the case $U(t,x)$ is symmetric and positive definite.
-Since $m$ already is determined by (3.4), we only need to prove that there is a unique symmetric solution $u(t,x)$ to equation (3.5).
-In that case, we obtain from Lemma A.1 a unique analytic antisymmetric $B(t,x)$ that solves
+By substituting $\bf m$} in $A$ we find that $A$ depends linearly on $A_{j}$, j=0, 1j=0,\,1, and that there exists an analytic and antisymmetric $B(t,x)$ so that $2U{\bf u}^{*}=A+B$, which gives $u^{*}=U^{-1}(A+B)/2$ and the solution
+(3.4) ${\bf u}(t,x)=\frac{1}{2}(A(t,x)-B(t,x))(U^{-1})^{*}(t,x)$
+This gives the surjectivity of $d\mathcal{F}$ in (3.3) and by the inverse function theorem we obtain local surjectivity of the nonlinear mapping $\mathcal{F}$ in (3.1) in a neighborhood $V$ of $(U,0)$.
+Observe that $\mathcal{F}$ will also be surjective on some $(F_{1},F_{0})$ such that $|F_{0}(t,0)|\ll 1$.
+For any symmetric ${\mathbb{M}}(t,x)$ such that $\partial_{t}^{m}{\mathbb{M}}(t,x)\equiv 0$ and ${\mathbb{M}}(t,0)=0$, there exists a neighborhood $\omega$ of $(0,0)$ such that $(U,{\mathbb{M}})\in V$ at $\omega$, and then (3.4) gives a solution to (3.1) in $\omega$.
+But (3.3) is not injective, since any antisymmetric $B$ gives a different solution.
+Actually, we have for fixed $U$ and ${\mathbb{M}}\equiv 0$ that the mapping $(A_{0},A_{1},B)\mapsto({\bf u},{\bf m})$ is bijective near $(0,0)$.
+It remains to prove uniqueness in the case $U(t,x)$ is positive definite near $(0,0)$, which we obtain when $U(0,0)=(F_{1}(0,0)/m!)^{1/2}>0$.
+Since $\bf m$} already is determined, we shall prove that there is only one symmetric solution ${\bf u}(t,x)$ to equation (3.4).
+If $\operatorname{Im}u^{*}=0$ then we obtain from Lemma A.1 a unique analytic antisymmetric $B(t,x)$ that solves
 $\operatorname{Im}(U^{-1}B)=-\operatorname{Im}(U^{-1}A)\qquad\text{near~$(0,0)$}$
 for the analytic $A(t,x)$ in (3.4). Thus the unique analytic symmetric solution is given by
-$u(t,x)=\frac{1}{2}U^{-1}(t,x)(A(t,x)+B(t,x))$
+${\bf u}(t,x)=\frac{1}{2}U^{-1}(t,x)(A(t,x)+B(t,x))$
 We find that $d\mathcal{F}$ is bijective which gives that $\mathcal{F}$ is a analytic bijection near $(0,0)$ in the analytic symmetric case. ∎
 4. Proof of Theorem 1.1 in the smooth case
-The smooth $N\times N$ systems form a local Fréchet spaces with the $C^{\infty}(\overline{\omega})$ norms, which are tame if $\overline{\omega}$ is a compact manifold with boundary, see Theorem C.8.
-For tame Fréchet spaces one can use the Nash–Moser theorem, which is the inverse function theorem for tame Fréchet spaces, see Theorem C.12.
-The difference to the analytic case is that is it not enought to check the local surjectivity of the linearized map when $M\equiv 0$, instead one has to check that for any $U$ and $M$ in a neighborhood.
-As before, it suffices to show that if $|U(0,0)|\neq 0$, $M(x)$ is symmetric and $M(0)=0$ then the following $C^{\infty}$ mapping is surjective near $(0,0)$
-(4.1) $C^{\infty}\ni(U(t,x),M(x))\overset{\mathcal{F}}{\longrightarrow}\\ \left(\partial_{t}\big(U(t,x)(t\operatorname{Id}_{N}+M(x))U^{*}(t,x)\big),U_{0}(x)M(x)U_{0}^{*}(x)\right)\\ =(F_{1}(t,x),F_{0}(x))\in C^{\infty}$
-where $U_{0}(x)=U({0,x})$.
-Here $F_{0}$ and $F_{1}$ are symmetric, $F_{1}(0,0)>0$ and $F_{0}(0)=0$.
-By (3.2) we then obtain $F$ from $F_{0}$ and $F_{1}$.
-It follows from Example C.4 that the mapping $\mathcal{F}$ is tame of degree 1.
-For any such $(F_{1},F_{0})$ we may take smooth $(U(t,x),M(x))$ so that $U(0,0)U^{*}(0,0)=F_{1}(0,0)>0$ and $M(0)=0$.
-Since we only need local solutions, we may assume that $U(t,x)\in C^{\infty}(\overline{\omega}_{n+1})$ so that $U(0,0)$ is invertible, $M(x)\in C^{\infty}(\overline{\omega}_{n})$ is symmetric and $M(0)=0$, where $\overline{\omega}_{k}\subset\mathbf{R}^{k}$ is a compact neighborhood of the origin.
-Let $V$ be a bounded neighborhood of $(U(t,x),M(x))$ in $C^{\infty}(\overline{\omega}_{n+1})\times C^{\infty}(\overline{\omega}_{n})$ so that the first component is invertible and the second component is symmetric.
-Since $M(0)=0$ we may for any $\varepsilon>0$ choose $V$ and a compact neighborhood $\overline{\omega}$ of the origin so that $(\overline{U}(t,x),\overline{M}(x))\in V$ gives $\|\overline{M}(x)\|\leq\varepsilon$ when $x\in\overline{\omega}$.
-For any $(U(t,x),M(x))\in V$ the differential of $\mathcal{F}$ at $(U(t,x),M(x))$ is equal to
-(4.2) $C^{\infty}\ni(u(t,x),m(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \left(\partial_{t}\big(2\operatorname{Re}(U(t\operatorname{Id}_{N}+M(x))u^{*})+UmU^{*}\big),U_{0}mU_{0}^{*}+2\operatorname{Re}(U_{0}Mu_{0}^{*})\right)\\ =(A_{1}(t,x),A_{0}(x))\in C^{\infty}$
-where $u_{0}(x)=u(0,x)$, and $m$, $A_{0}$ and $A_{1}$ are symmetric.
-We shall show that if $\overline{\omega}_{n+1}$ and $\overline{\omega}_{n}$ are sufficiently small, then the mapping $d\mathcal{F}$ is tame and surjective with a tame right inverse, which will show the surjectivity of $\mathcal{F}$ by Theorem C.12.
-For that, we need that $\|M(x)\|\ll 1$ for $|x|\ll 1$.
-We find from (4.2) that
-(4.3) $m=U_{0}^{-1}\Big(A_{0}-2\operatorname{Re}(U_{0}Mu^{*}_{0})\Big)(U_{0}^{-1})^{*}$
-is linear in $Mu_{0}^{*}$ and if $P(t,M(x))=t\operatorname{Id}_{N}+M(x)$ then
-(4.4) $2\operatorname{Re}UP(t,M(x))u^{*}=\int_{0}^{t}A_{1}(s,x)\,ds-U(t,x)m(x)U^{*}(t,x)+R(x)=A(t,x)$
-for some smooth symmetric $R(x)$. Observe that $A$ depends linearly on $m$ and thus linearly on $Mu_{0}^{*}$ by (4.3). By (4.4) there exists a smooth antisymmetric $B(t,x)$ so that
-(4.5) $2P(t,M(x))u^{*}(t,x)=\\ U^{-1}(t,x)\int_{0}^{t}A_{1}(s,x)\,ds-m(x)U^{*}(t,x)+U^{-1}(t,x)R(x)+U^{-1}(t,x)B(t,x)$
-Since $R(x)$ is symmetric and $B(0,x)$ is antisymmetric, we may replace $R(x)$ with $R(x)+B(0,x)$ and let $B(0,x)\equiv 0$. Then we can replace $B(t,x)$ with $tB(t,x)$, which gives
-(4.6) $U^{-1}(t,x)tB(t,x)=P(t,M(x))U^{-1}(t,x)B(t,x)-M(x)U^{-1}(t,x)B(t,x)$
-Now we shall use the symmetric Malgrange division theorem given by Corollary B.4 on every term on the right hand side of (4.5) using (4.6).
-That will give a $C^{\infty}$ surjective tame map with a $C^{\infty}$ tame right inverse by Example C.13.
-First, if $P=P(t,M(x))$ then Corollary B.4 gives smooth $C_{1}(t,x)$ and $R_{1}(x)$ so that
-(4.7) $U^{-1}(t,x)\int_{0}^{t}A_{1}(s,x)\,ds=PC_{1}(t,x)+R_{1}(x)\qquad\text{near $(0,0)$}$
-where $C_{1}$ and $R_{1}$ depend linearly on $m$ and thus linearly on $Mu_{0}^{*}$ by (4.3), and that
-(4.8) $m(x)U^{*}(t,x)=PC_{2}(t,x)+R_{2}(x)\qquad\text{near $(0,0)$}$
-where $C_{2}$ and $R_{2}$ are smooth and depend linearly on $Mu_{0}^{*}$. We also find
-(4.9) $U^{-1}(t,x)=PC_{3}(t,x)+R_{3}(x)\qquad\text{near $(0,0)$}$
-with smooth $C_{3}$ and $R_{3}$ such that $R_{3}(0)=U^{-1}(0,0)$ is invertible, thus $R_{3}(x)$ is invertible when $|x|\ll 1$. Finally, we obtain smooth $C_{4}(t,x)$ and $R_{4}(x)$ so that
-(4.10) $M(x)U^{-1}(t,x)B(t,x)=PC_{4}(t,x)+R_{4}(x)\qquad\text{near $(0,0)$}$
-and $C_{4}(t,0)=R_{4}(0)=0$.
-Here $C_{4}$ and $R_{4}$ depend linearly on $B$ but the differentials with respect to $B$ is $\mathcal{O}(M)$.
-We have that $C_{j}$ and $R_{j}$ depend tamely on $U$ and $M$, $\forall\,j$.
-Summing up, we find from (4.6)–(4.10) that
-(4.11) $2Pu^{*}(t,x)=P(C_{1}+C_{2}+C_{3}R+C_{4}+U^{-1}B)+R_{1}+R_{2}+R_{3}R+R_{4}=PC_{0}+R_{0}$
-near $(0,0)$, where $C_{0}(t,x)$ and $R_{0}(x)$ are smooth depending linearly on $R$ $Mu_{0}^{*}$ and $B$. Thus for fixed $U$, $M$ and $B$, we can write
-(4.12) $\left\{\begin{aligned} &C_{0}=C_{00}+C_{01}Mu_{0}^{*}+C_{02}R+C_{03}B+U^{-1}B\\ &R_{0}=R_{00}+R_{01}Mu_{0}^{*}+R_{3}R+R_{03}B\end{aligned}\right$.
-where $R_{3}$ and $U^{-1}$ are invertible.
-Observe that, except $U^{-1}B$ and $R_{3}R$, the products are not matrix multiplications, instead they represent linear maps.
-We have that the norms $\|C_{03}\|$ and $\|R_{03}\|$ are bounded by $C\|M\|$.
-Now, since $u_{0}(x)$ appears in the equation we shall let $G(t,x)=G_{0}(x)+tG_{1}(t,x)$ for any smooth $G$, and $C_{jk}(t,x)=C_{jk}^{0}(x)+tC_{jk}^{1}(t,x)$, $\forall\,jk$.
-For fixed $U$, $M$ and $B$ we obtain that $2u^{*}\equiv C_{0}$ with $R_{0}\equiv 0$ if
-(4.13) $\left\{\begin{aligned} &2u_{0}^{*}-C_{01}^{0}Mu_{0}^{*}-C_{02}^{0}R\equiv C_{00}^{0}+(C_{03}B)_{0}+(U^{-1}B)_{0}\\ &-tC_{01}^{1}Mu_{0}^{*}+2tu_{1}^{*}-tC_{02}^{1}R\equiv tC_{00}^{1}+t(C_{03}B)_{1}+t(U^{-1}B)_{1}\\ &R_{01}Mu_{0}^{*}+R_{3}R\equiv-R_{00}-R_{03}B\end{aligned}\right$.
-By dividing the second equation by $t$, this system has the following matrix for $(u^{*}_{0},u_{1}^{*},R)$
-(4.14) $\begin{pmatrix}2\operatorname{Id}_{N}-C_{01}^{0}M&0&-C_{02}^{0}\\ -C_{01}^{1}M&2\operatorname{Id}_{N}&-C_{02}^{1}\\ R_{01}M&0&R_{3}\end{pmatrix}$
-which is invertible when $M\equiv 0$ since $R_{3}$ is invertible, and therefore when $\|M(x)\|\ll 1$.
-Thus, if the neighborhood $V$ is small enough, then for any $B$ we obtain a unique solution $u^{*}\equiv u_{0}^{*}+tu_{1}^{*}\equiv C_{0}/2$ with $R_{0}\equiv 0$ in a sufficiently small compact neighborhood $\overline{\omega}$ of $(0,0)$.
-This gives a smooth tame map by Example C.4, and the neighborhood $\overline{\omega}$ only depends on the bounds on $U$, $M$ and $B$.
-Having found $u^{*}$ we obtain $(u,m)$ by (4.3) in $\overline{\omega}$.
-This gives surjectivity and a tame right inverse to $d\mathcal{F}$ locally for any $(U,M)\in V$.
-By Theorem C.12, we obtain that the $C^{\infty}$ mapping $\mathcal{F}$ has a tame local right inverse on a neighborhood of $(F_{1},F_{0})$ in (4.1), giving a solution to (1.6).
-Observe that $\mathcal{F}$ is surjective also on some $(F_{1},F_{0})$ such that $F_{0}(0)\neq 0$ depending on $F_{1}$.
-Note that every antisymmetric $B$ gives a solution.
-It remains to prove uniqueness in the symmetric case for $(U,M)\in V$ when $U(t,x)>0$ and $u(t,x)$ is symmetric.
-We shall then show that there is only one symmetric solution to (1.6), by finding the corresponding $B(t,x)$ in (4.5).
-By adding the equation $\operatorname{Im}u^{*}\equiv\operatorname{Im}(C_{0}/2)\equiv 0$ with $C_{0}$ given by (4.12), we find
-(4.15) $\left\{\begin{aligned} &2u_{0}^{*}-C_{01}^{0}Mu_{0}^{*}-C_{02}^{0}R-(C_{03}B)_{0}-(U^{-1}B)_{0}\equiv C_{00}^{0}\\ &-tC_{01}^{1}Mu_{0}^{*}+2tu_{1}^{*}-tC_{02}^{1}R-t(C_{03}B)_{1}-t(U^{-1}B)_{1}\equiv tC_{00}^{1}\\ &R_{01}Mu^{*}+R_{3}R+R_{03}B\equiv-R_{00}\\ &\operatorname{Im}(C_{01}Mu_{0}^{*})+\operatorname{Im}(C_{02}R)+\operatorname{Im}(C_{03}B)+\operatorname{Im}(U^{-1}B)\equiv-\operatorname{Im}(C_{00})\end{aligned}\right$.
-By dividing the second equation by $t$, this system has following matrix for $(u_{0}^{*},u_{1}^{*},R,B)$
-(4.16) $\begin{pmatrix}2\operatorname{Id}_{N}-C_{01}^{0}M&0&-C_{02}^{0}&-(C_{03}\,\boldsymbol{\cdot})_{0}-(U^{-1}\,\boldsymbol{\cdot})_{0}\\ -C_{01}^{1}M&2\operatorname{Id}_{N}&-C_{02}^{1}&-(C_{03}\,\boldsymbol{\cdot})_{1}-(U^{-1}\,\boldsymbol{\cdot})_{1}\\ R_{01}M&0&R_{3}&R_{03}\\ \operatorname{Im}(C_{01}M\,\boldsymbol{\cdot})&0&\operatorname{Im}(C_{02}\,\boldsymbol{\cdot})&\operatorname{Im}(C_{03}\,\boldsymbol{\cdot})+\operatorname{Im}(U^{-1}\,\boldsymbol{\cdot})\end{pmatrix}$
-where $C\,\boldsymbol{\cdot}$ is the map $E\to CE$.
-Thus $\operatorname{Im}(U^{-1}\,\boldsymbol{\cdot})$ is the map $E\mapsto\operatorname{Im}(U^{-1}E)$, which is invertible by Lemma A.1 because $U^{-1}>0$.
-Since $R_{3}$ also is invertible and $C_{03}=R_{03}=0$ when $M\equiv 0$, we find that (4.16) is invertible when $M\equiv 0$, and therefore when $\|M(x)\|\ll 1$.
-Thus, if the neighborhood $V$ is small enough, then we obtain a unique smooth solution $(u_{0}^{*},u_{1}^{*},R,B)$ to (4.15) near $(0,0)$ giving symmetric $u\equiv u_{0}+tu_{1}\equiv C_{0}/2$ with $R_{0}\equiv 0$.
-By (4.3) this gives a unique symmetric solution $(u,m)\in C^{\infty}$ to (4.2).
-Observe that this gives a tame map by Example C.4, and the neighborhood $\overline{\omega}$ only depends on the bounds on $(U,M)\in V$.
-Thus we obtain in the positive definite case that $d\mathcal{F}$ is a tame bijection locally for any positive definite $U$ and symmetric $M$ such that $(U,M)\in V$, so by Theorem C.12 we obtain that the mapping $\mathcal{F}$ is a tame bijection locally in this case, giving a unique solution to (1.6).
+The smooth $N\times N$ systems form a local Fréchet spaces with the $C^{\infty}(\overline{\omega})$ norms, which are tame if $\overline{\omega}$ is a compact neighborhood with boundary, see Theorem B.8.
+For tame Fréchet spaces one can use the Nash–Moser theorem, which is the inverse function theorem for tame Fréchet spaces.
+The difference to the analytic case is that is it not enought to check the local surjectivity of the linearized map when ${\mathbb{M}}(0)\equiv 0$, instead one has to check that for any smooth $U$ and $\mathbb{M}$} in a neighborhood, see Theorem B.12.
+So we shall show that if $U(t,x)\in C^{\infty}$, $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)\in C^{\infty}$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, then the following $C^{\infty}$ mapping is surjective near $(0,0)$
+(4.1) $C^{\infty}\ni(U(t,x),{\mathbb{M}}(t,x))\overset{\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(U(t,x)(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x))U^{*}(t,x)\big),T_{m}(U{\mathbb{M}}U^{*})(t,x))\Big)\\ =(F_{1}(t,x),F_{0}(t,x))\in C^{\infty}$
+Here $F_{j}$ is symmetric, j=0, 1,j=0,\,1, $F_{1}(0,0)>0$, $\partial_{t}^{m}F_{0}\equiv 0$ and $F_{0}(t,0)\equiv 0$.
+As before, if we obtain a solution $(F_{1},F_{0})$ to (4.1) then we obtain $F(t,x)$ from (3.2) so that $F_{0}(t,x)=T_{m}(F)(t,x)$ and $F_{1}(t,x)=\partial_{t}^{m}F(t,x)$.
+It follows from Example B.4 that the mapping $\mathcal{F}$ is tame of degree $m$.
+Now, for any such $(F_{1}(t,x),F_{0}(t,x))$ we may choose $(U_{0}(t,x),{\mathbb{M}}_{0}(t,x))$ so that $U_{0}(t,x)$ is invertible and $U_{0}(0,0)U^{*}_{0}(0,0)\equiv F_{1}(0,0)$ and ${\mathbb{M}}_{0}(t,x)$ is a symmetric polynomial in $t$ of degree less than $m$ such that ${\mathbb{M}}_{0}(t,0)\equiv 0$.
+Observe that we may choose symmetric $U_{0}$ such that $U_{0}(0,0)=F_{1}^{1/2}(0,0)>0$, and then $U_{0}>0$ near the origin.
+Since we shall only need local solutions, we may assume that $U_{0}(t,x)$ and ${\mathbb{M}}_{0}(t,x)\in C^{\infty}(\overline{\omega})$ where $\overline{\omega}$ is compact neighborhood of the origin so that $U(t,x)$ is invertible and ${\mathbb{M}}(t,x)$ is symmetric such that $\partial_{t}^{m}{\mathbb{M}}(t,x)\equiv 0$ and ${\mathbb{M}}(t,0)\equiv 0$.
+Then we let $V$ be a bounded neighborhood of $(U_{0}(t,x),{\mathbb{M}}_{0}(t,x))$ in $C^{\infty}(\overline{\omega})$ so that the first component is invertible and the second component is symmetric polynomial in $t$ of degree less than $m$.
+If $U_{0}>0$ then $V$ is a neighborhood with symmetric matrices.
+Since ${\mathbb{M}}_{0}(t,0)\equiv 0$ we may for $\varepsilon>0$ choose $V$ and a compact neighborhood $\overline{\omega}$ so that $(U,{\mathbb{M}})\in V$ gives $\|{\mathbb{M}}\|\leq\varepsilon$ in $\overline{\omega}$.
+We find for any $(U(t,x),{\mathbb{M}}(t,x))\in V$ that the differential of $\mathcal{F}$ at $(U(t,x),{\mathbb{M}}(t,x))$ is equal to
+(4.2) $C^{\infty}\ni({\bf u}(t,x),{\bf m}(t,x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(2\operatorname{Re}(U(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x)){\bf u}^{*})+U{\bf m}U^{*}\big),T_{m}(U{\bf m}U^{*})+2\operatorname{Re}T_{m}(U{\mathbb{M}}{\bf u}^{*})\Big)\\ =(A_{1}(t,x),A_{0}(t,x))\in C^{\infty}$
+where ${\bf u}(t,x)$ and ${\bf m}(t,x)$ are smooth such that $\bf m$} is symmetric and $\partial_{t}^{m}{\bf m}\equiv 0$ which gives smooth symmetric $A_{j}$, j=0, 1j=0,\,1, such that $\partial_{t}^{m}A_{0}\equiv 0$.
+Next, we shall consider the special case when the functions are constant in $x$.
+We shall show that if ${\bf u}(t)$ and ${\bf m}(t)\in C^{\infty}(\overline{\omega})$ for a sufficiently small neighborhoods $\overline{\omega}$ of the origin in $\mathbf{R}$, then the mapping $d\mathcal{F}$ is tame and surjective with a tame right inverse.
+For that, the condition that $\|{\mathbb{M}}\|\ll 1$ in $\overline{\omega}$ is necessary.
+Proposition 4.1.
+Let $U(t)$ and ${\mathbb{M}}(t)$ be smooth and $N\times N$ valued functions such that $|U(t)|\neq 0$, $\partial^{m}_{t}{\mathbb{M}}(t)\equiv 0$ and $\|{\mathbb{M}}(t)\|\ll 1$.
+Then in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin, the mapping $d\mathcal{F}$ given by (4.2) is surjective, having a tame local right inverse of degree 3 in $\overline{\omega}$.
+If $U(t)$ is positive definite, then $d\mathcal{F}$ is bijective and has a tame local inverse of degree 3.
+By Theorem B.12, we obtain that the $C^{\infty}$ tame mapping $\mathcal{F}$ is surjective on a neighborhood $W$ of a given $(F_{1}(t),F_{0}(t))$ close to the origin, having a smooth tame right inverse $\mathcal{F}^{-1}$ giving $(U(t),{\mathbb{M}}(t))=\mathcal{F}^{-1}(F_{1}(t),F_{0}(t))$.
+Now, if $F_{j}(t)=F_{j}(t,0)$ then for $|x|\ll 1$ we obtain that $(F_{1}(t,x),F_{0}(t,x))\in W$, so we obtain the symmetric preparation (1.7) with
+(4.3) $(U(t,x),{\mathbb{M}}(t,x))=\mathcal{F}^{-1}(F_{1}(t,x),F_{0}(t,x))$
+in a neighborhood of $(0,0)$.
+If $U(t)$ is positive definite, then by Theorem B.12 we obtain that the mapping $\mathcal{F}$ is a bijection in a neighborhood $W$ of a given $(F_{1}(t),F_{0}(t))$ near the origin with the tame inverse $\mathcal{F}^{-1}$ giving a unique $(U(t),{\mathbb{M}}(t))=\mathcal{F}^{-1}(F_{1}(t),F_{0}(t))$.
+If $F_{j}(t)=F_{j}(t,0)$ and $|x|\ll 1$ then we obtain from (4.3) the unique symmetric solution $(U(t,x),{\mathbb{M}}(t,x))$ to (1.7) near the origin which finishes the proof of Theorem 1.1 in the smooth case.
+∎
+It remains to prove Proposition 4.1. We find from (4.2) that
+(4.4) ${\bf m}=T_{m}\left(U^{-1}\big(A_{0}-2\operatorname{Re}(U{\mathbb{M}}{\bf u}^{*})\big)(U^{-1})^{*}\right)$
+and if $P(t,{\mathbb{M}}(t))=t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t)$ with symmetric ${\mathbb{M}}(t)=\sum_{j<m}t^{j}M_{j}$ we have
+(4.5) $2\operatorname{Re}UP(t,{\mathbb{M}}(t)){\bf u}^{*}=I_{m}(A_{1})(t)-U(t){\bf m}(t)U^{*}(t)+\mathbb{R}(t)=A(t)$
+for some smooth symmetric $\mathbb{R}(t)$ such that $\partial_{t}^{m}\mathbb{R}(t)\equiv 0$.
+Observe that $A$ depends linearly on $\bf m$} by (4.4) and thus linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$, and that we may assume that $\|{\mathbb{M}}\|=\sum_{j<m}\|M_{j}\|\ll 1$.
+By (4.5) there exists a smooth antisymmetric $B(t)$ so that
+(4.6) $2P(t,{\mathbb{M}}(t)){\bf u}^{*}(t)=U^{-1}(t)I_{m}(A_{1})(t)-{\bf m}(t)U^{*}(t)+U^{-1}(t)\mathbb{R}(t)+U^{-1}(t)B(t)$
+We shall keep $B$ fixed in the following, and we may take $B(t)\equiv 0$.
+Since $\mathbb{R}(t)$ is symmetric and $B(t)$ is antisymmetric, we may replace $\mathbb{R}(t)$ with $\mathbb{R}(t)-T_{m}(B)(t)$ and let $T_{m}(B)\equiv 0$.
+Then we can replace $B(t)$ by $t^{m}B(t)$ which gives
+(4.7) $U^{-1}(t)t^{m}B(t)=P(t,{\mathbb{M}}(t))U^{-1}(t)B(t)-{\mathbb{M}}(t)U^{-1}(t)B(t)$
+We shall use the following result, which is the adjoint version of Proposition 3.2 in [3] in the special case when $\pi_{m}=\operatorname{Id}_{N}$.
+Proposition 4.2.
+Let $G(t)\in\mathcal{S}(\mathbf{R})$ be $N\times N$ valued and let $P(t,{\mathbb{M}}(t))=t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t)$ with $N\times N$ system ${\mathbb{M}}(t)=\sum_{j<m}t^{j}M_{j}$, ${\mathbb{M}}=\left\{\,M_{j}\,\right\}_{j<m}$ and $\|{\mathbb{M}}\|=\sum_{j<m}\|M_{j}\|$.
+Then we can find $N\times N$ systems $Q(t,{\mathbb{M}},G)\in C^{\infty}$ and $\mathbb{R}(t,{\mathbb{M}},G)=\sum_{j<m}t^{j}R_{j}({\mathbb{M}},G)\in C^{\infty}$ depending linearly on $G(t)$, such that
+(4.8) $G(t)=P(t,{\mathbb{M}}(t))Q(t,{\mathbb{M}},G)+\mathbb{R}(t,{\mathbb{M}},G){\quad}\text{if}{\quad}\|{\mathbb{M}}\|<1$
+Also, we have the following estimates when $\|{\mathbb{M}}\|<1$
+(4.9) $\|\partial_{t}^{{\alpha}}\partial_{{\mathbb{M}}}^{{\beta}}Q(t,{\mathbb{M}},G)\|\leq C_{{\alpha}{\beta}}\int(\|G\|+\|G^{(j)}\|)\,dt{\quad}j={\alpha}+3+m(|{\beta}|+1)$
+$\|\partial_{{\mathbb{M}}}^{{\beta}}R_{k}(t,{\mathbb{M}},G)\|\leq C_{{\beta}}\int(\|G\|+\|G^{(j)}\|)\,dt{\quad}j=2+m(|{\beta}|+1)\qquad k<m$
+We find that (4.8) gives a smooth surjective tame map $C^{\infty}\ni(Q,\mathbb{R})\mapsto G\in\mathcal{S}$ of degree 0 having a smooth tame right inverse of degree $3+m$, see Example B.14.
+Observe that the estimate (4.8) is not tame in the parameters $\mathbb{M}$}.
+Thus, for $\|{\mathbb{M}}\|\ll 1$ we shall use Proposition 4.2 on every term on the right hand side of (4.6) using (4.7), after cut off so that they have compact support.
+We will denote by $\mathbb{R}_{k}(t)$ the terms that are polynomials of degree less than $m$ in $t$.
+First, if $P=P(t,{\mathbb{M}}(t))$ then Proposition 4.2 gives smooth $C_{1}(t)$ and $\mathbb{R}_{1}(t)$ so that
+(4.10) $U^{-1}(t)I_{m}(A_{1})(t)=PC_{1}(t)+\mathbb{R}_{1}(t)\qquad\text{near $t=0$}$
+where $C_{1}$ and $\mathbb{R}_{1}$ are smooth depending linearly on $\bf m$} and thus linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$ by (4.4), and that
+(4.11) ${\bf m}(t)U^{*}(t)=PC_{2}(t)+\mathbb{R}_{2}(t)\qquad\text{near $t=0$}$
+where $C_{2}$ and $\mathbb{R}_{2}$ are smooth and depending linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$. We also find
+(4.12) $U^{-1}(t)=PC_{3}(t)+\mathbb{R}_{3}(t)\qquad\text{near $t=0$}$
+with smooth $C_{3}$ and $\mathbb{R}_{3}$, thus $U^{-1}\mathbb{R}=PC_{3}\mathbb{R}+\mathbb{R}_{3}\mathbb{R}$.
+By (4.9), $C_{3}$ is uniformly bounded when $\|{\mathbb{M}}\|<1$.
+Thus $\mathbb{R}_{3}(0)$ is invertible if $\|{\mathbb{M}}(0)\|\ll 1$ since $\mathbb{R}_{3}(0)=U^{-1}(0)-{\mathbb{M}}(0)C_{3}(0)$.
+Finally, we obtain smooth $C_{4}(t)$ and $\mathbb{R}_{4}(t)$ so that
+[…]
+where $\|C_{4}\|_{k}$ and $\|\mathbb{R}_{4}\|_{k}$ are $\mathcal{O}(\|{\mathbb{M}}\|)$ for any local $C^{k}$ norm.
+We find from Proposition 4.2 that $C_{j}(t)$ and $\mathbb{R}_{j}(t)$, $\forall\,j$, depend smoothly and tamely of degree $3+m$ on $U(t)$ and ${\mathbb{M}}(t)$.
+Summing up, we find from (4.7), and (4.10)–(4.13) that
+(4.14) $2Pu^{*}(t)=P\left(C_{1}+C_{2}+C_{3}\mathbb{R}+C_{4}B+U^{-1}B\right)+\mathbb{R}_{1}+\mathbb{R}_{2}+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B=PC_{0}+\mathbb{R}_{0}$
+locally near the origin in $\mathbb{R}$.
+Here $C_{0}$ and $\mathbb{R}_{0}$ are smooth depending linearly on $\mathbb{R}$, $B$ and $T_{m}({\mathbb{M}}{\bf u}^{*})$.
+Thus for fixed $U$ and $\mathbb{M}$} we can write
+(4.15) $\left\{\begin{aligned} &C_{0}=C_{00}+C_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+C_{3}\mathbb{R}+C_{4}B+U^{-1}B\\ &\mathbb{R}_{0}=\mathbb{R}_{00}+\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B\end{aligned}\right$.
+where $\mathbb{R}_{3}(t)$ is invertible near $t=0$ and $\|C_{4}\|_{k}+\|R_{4}\|_{k}=\mathcal{O}(\|{\mathbb{M}}\|)$, $\forall\,k$.
+Observe that the products are matrix multiplications, except $C_{01}T_{m}({\mathbb{M}}{\bf u}^{*})$ and $R_{01}T_{m}({\mathbb{M}}{\bf u}^{*})$, which represent linear maps.
+Now, since $T_{m}({\mathbb{M}}{\bf u}^{*})$ appears in the equations we shall put $G(t)=T_{m}(G)(t)+t^{m}G_{1}(t)$ for $G\in C^{\infty}$ and $C_{jk}(t)=T_{m}(C_{jk})(t)+t^{m}C_{jk}^{1}(t)$, $\forall\,jk$.
+Then for fixed $U$, $\mathbb{M}$} and $B$ we obtain that $2{\bf u}^{*}\equiv C_{0}$ and $\mathbb{R}_{0}\equiv 0$ if
+(4.16) $\left\{\begin{aligned} &2T_{m}({\bf u}^{*})-T_{m}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))-T_{m}(C_{3}\mathbb{R})\equiv T_{m}(C_{00})+T_{m}(C_{4}B)+T_{m}(U^{-1}B)\\ &-C_{01}^{1}T_{m}({\mathbb{M}}{\bf u}^{*})+2{\bf u}_{1}^{*}-(C_{3}\mathbb{R})_{1}\equiv C_{00}^{1}+(C_{4}B)_{1}+(U^{-1}B)_{1}\\ &\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}\equiv-\mathbb{R}_{00}-\mathbb{R}_{4}B\end{aligned}\right$.
+Observe that the first equation is modulo $\mathcal{O}(t^{m})$, and that
+(4.17) $\|T_{m}({\mathbb{M}}{\bf u}^{*})\|_{k}\lesssim\|{\mathbb{M}}\|\|{\bf u}^{*}\|_{k+m-1}$
+since $\|{\mathbb{M}}\|_{k}\lesssim\|{\mathbb{M}}\|$ and $\|T_{m}{\bf u}^{*}\|_{k}\lesssim\|{\bf u}^{*}\|_{k+m-1}$ for any local $C^{k}$ norm. Then we get from (4.17) that (4.18) is given by
+(4.18) $\begin{pmatrix}2\operatorname{Id}_{N}&0&-T_{m}(C_{3}\ \boldsymbol{\cdot})\\ 0&2\operatorname{Id}_{N}&-(C_{3}\ \boldsymbol{\cdot})_{1}\\ 0&0&\mathbb{R}_{3}\end{pmatrix}$
+modulo terms having local $C^{k}$ norms that are $\mathcal{O}(\|{\mathbb{M}}\|)$.
+Here $\operatorname{Im}(C\ \boldsymbol{\cdot})$ is the mapping $E\mapsto\operatorname{Im}(CE)$ and $T_{m}(C\ \boldsymbol{\cdot})$ is the mapping $E\mapsto T_{m}(CE)$.
+The system (4.18) is invertible close to $t=0$ since $\mathbb{R}_{3}(0)$ is invertible.
+Thus, we find that the system (4.16) is invertible in a compact neighborhood $\overline{\omega}$ of the origin if $\|{\mathbb{M}}\|\ll 1$, which can be obtained by shrinking $V$ and $\overline{\omega}$.
+Since this system is linear, smooth and invertible, we find from Example B.4 that it has a tame smooth inverse of degree 0 which gives a solution $(T_{m}({\bf u}^{*}),{\bf u}^{*}_{1},\mathbb{R})$.
+Thus, we obtain a smooth solution ${\bf u}^{*}(t)\equiv T_{m}({\bf u}^{*})(t)+t^{m}{\bf u}^{*}_{1}(t)\equiv C_{0}(t)/2$ for $\mathbb{R}_{0}(t)\equiv 0$ in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin.
+Observe that the neighborhood $\overline{\omega}$ only depends on the bounds on $U$, $\mathbb{M}$} and $B$.
+Having found ${\bf u}^{*}(t)$ satisfying (4.5) we obtain $({\bf u}(t),{\bf m}(t))$ by (4.4) in $\overline{\omega}$.
+This gives surjectivity and a tame local right inverse of degree $3+m$ to $d\mathcal{F}$ in $\overline{\omega}$ for any $(U(t),{\mathbb{M}}(t))\in V$.
+Observe that different antisymmetric $B$ gives different inverses.
+It remains to prove uniqueness in the symmetric case for $(U,{\mathbb{M}})\in V$, when $U$ and ${\bf u}(t)$ are symmetric and $U>0$ which then holds in $V$.
+We shall show that there is a unique symmetric solution to the equation, and find the corresponding $B(t)$ in (4.6).
+By adding the equation $\operatorname{Im}{\bf u}(t)\equiv\operatorname{Im}(C_{0})(t)\equiv 0$ with $C_{0}(t)$ given by (4.15) we find as before
+(4.19) $\left\{\begin{aligned} &2T_{m}({\bf u})-T_{m}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))-T_{m}(C_{3}\mathbb{R})-T_{m}(C_{4}B)+T_{m}(U^{-1}B)\equiv T_{m}(C_{00})\\ &-C_{01}^{1}T_{m}({\mathbb{M}}{\bf u}^{*})+2{\bf u}_{1}-(C_{3}\mathbb{R})_{1}-(C_{4}B)_{1}-(U^{-1}B)_{1}\equiv C_{00}^{1}\\ &\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B\equiv-\mathbb{R}_{00}\\ &\operatorname{Im}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))+\operatorname{Im}(C_{3}\mathbb{R})+\operatorname{Im}(C_{4}B)+\operatorname{Im}(U^{-1}B)\equiv-\operatorname{Im}(C_{00})\end{aligned}\right$.
+which is invertible near $t=0$ when $\|{\mathbb{M}}\|\ll 1$. In fact, the matrix of the system is
+(4.20) $\begin{pmatrix}2\operatorname{Id}_{N}&0&-T_{m}(C_{3}\ \boldsymbol{\cdot})&-T_{m}(U^{-1}\ \boldsymbol{\cdot})\\ 0&2\operatorname{Id}_{N}&-(C_{3}\ \boldsymbol{\cdot})_{1}&-(U^{-1}\ \boldsymbol{\cdot})_{1}\\ 0&0&\mathbb{R}_{3}&0\\ 0&0&\operatorname{Im}(C_{3}\ \boldsymbol{\cdot})&\operatorname{Im}(U^{-1}\ \boldsymbol{\cdot})\end{pmatrix}$
+modulo terms having local $C^{k}$ norms that are $\mathcal{O}(\|{\mathbb{M}}\|)$.
+Here $\operatorname{Im}(U^{-1}\ \boldsymbol{\cdot})$ is invertible on antisymmetric matrices by Lemma A.1.
+Since $\mathbb{R}_{3}(0)$ also is invertible, we find that (4.20) is invertible near the origin.
+As before, we find that (4.19) is invertible in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin if $\|{\mathbb{M}}\|\ll 1$, which can be obtained by shrinking $V$ and $\overline{\omega}$.
+Thus, we obtain a unique smooth solution $({\bf u}(t),\mathbb{R}(t),B(t))$ to (4.19) with symmetric $\bf u$}, and by (4.4) a unique smooth symmetric solution $({\bf u}(t),{\bf m}(t))$ to (4.2).
+Thus, $d\mathcal{F}$ is a tame bijection in a compact neighborhood $\overline{\omega}$ of the origin for any symmetric $(U(t),{\mathbb{M}}(t))\in V$.
+Observe that this gives a tame inverse $d\mathcal{F}^{-1}$ of degree $3+m$ by Example B.14, and the neighborhood $\overline{\omega}$ only depends on the bounds on $U$ and $\mathbb{M}$}.
 ∎
@@
 To simplify, we define the symmetric product $S(U,A)=\frac{1}{2}(UA+AU)$, which is symmetric when $A$ is symmetric and antisymmetric when $A$ is antisymmetric, thus $A\mapsto S(U,A)$ preserves the symmetry.
-Since any $N\times N$ matrix can be written uniquely as a sum of a symmetric and an antisymmetric matrix, we find that $A\to S(U,A)$ is a bijection for any $N\times N$ matrix $A$.
 We also have that linear ON mappings $O$ preserves the symmetry, since $OAO^{-1}=OAO^{*}$.
+Next, we are going to prove that $A\mapsto S(U,A)$ is injective for both symmmetric and antisymmetric $A$.
+Since any $N\times N$ matrix can be written uniquely as a sum of a symmetric and an antisymmetric matrix, we will then find that $A\to S(U,A)$ is a bijection for any $N\times N$ matrix $A$.
 Thus, by using an ON base of eigenvectors to $U$, we find that $U=\{\mu_{j}\delta_{jk}\}_{jk}$ with $\mu_{j}>0,\ \forall\,j$, and $A=\{\alpha_{jk}\}_{jk}$.
-If $A$ is symmetric then $\alpha_{kj}=\overline{\alpha}_{jk}$ $\forall\,jk$, and if $A$ is antisymmetric then $\alpha_{kj}=-\overline{\alpha}_{jk}$ $\forall\,jk$.
+If $A$ is symmetric then $\alpha_{kj}=\alpha_{jk}$ $\forall\,jk$, and if $A$ is antisymmetric then $\alpha_{kj}=-\alpha_{jk}$ $\forall\,jk$.
 In both cases, we find that
@@
 ∎
-Appendix B Symmetric Malgrange division
-In this section, we shall prove a symmetric version of the Malgrange preparation theorem for systems, by modifying the proofs in the Appendix of [2] which generalized [5, Sect.
-7.5] to systems.
-In what follows, we shall let $|A|=\det A$ be the determinant of $N\times N$ matrices $A$, and let $\|A\|$ be the matrix norm.
-We are going to divide matrix valued analytic functions with symmetric matrix valued (first order) polynomials.
-We shall let $P(t,B)=t\operatorname{Id}_{N}+B$, where $B$ is symmetric $N\times N$ matrix, $\omega$} be an open set in $\mathbf{C}$ such that $|P(t,B)|\neq 0$ when $t\in\partial{\omega}\in C^{1}$.
-If $G(t)$ is analytic $N\times N$ system in $\overline{\omega}$, then we have
-(B.1) $G(t)=Q(t)P(t,B)+R{\qquad}t\in{\omega}$
-where
-(B.2) $Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds{\qquad}t\in{\omega}$
-is analytic and
-(B.3) $R$ $=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(P(s,B)-P(t,B))(s-t)^{-1}\,ds$
-$=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}\,ds$
-is constant.
-Observe that $Q(t)$ and $R$ depend analytically on $B$.
-The remainder RR\/ is uniquely determined by the condition that $RP(t,B)^{-1}$ is analytic when $t\not\in{\omega}$.
-In fact, it is easy to see that (B.3) gives
-$RP(t,B)^{-1}=-(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds{\qquad}t\not\in{\omega}$
-which is analytic in $\complement{\omega}$. Then the uniqueness of RR\/ follows, since (B.1) implies that
-$Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,B)^{-1}(s-t)^{-1}\,ds-(2{\pi}i)^{-1}\int_{\partial{\omega}}RP(s,B)^{-1}(s-t)^{-1}\,ds$
-when $t\in{\omega}$.
-If $RP(s,B)^{-1}$ is analytic in $\complement{\omega}$ we may push the integration contour to infinity in the second integral.
-Since the integrand is $O(|s|^{-2})$ in the first integral we obtain (B.2).
-Now we want to make this division in a small complex strip around $\mathbf{R}$, with uniform bounds.
-Lemma B.1.
-Let $G(t)$ be a bounded analytic $N\times N$ system in $\{\,t\in\mathbf{C}:|\operatorname{Im}t|<{\varepsilon}\,\}$.
-Then for any symmetric $N\times N$ matrix $B$ we can find analytic $N\times N$ system $Q^{{\varepsilon}}(t,B,G)$ and $R^{{\varepsilon}}(B,G)$ depending linearly on $G(t)$, such that
-(B.4) $G(t)=Q^{{\varepsilon}}(t,B,G)P(t,B)+R^{{\varepsilon}}(B,G){\quad}\text{if}{\quad}\|B\|<1$
-We also have the following estimates
-(B.5) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q^{{\varepsilon}}(t,B,G)\|\leq C_{{\alpha}{\beta}}M{\varepsilon}^{-{\alpha}-|{\beta}|-2}$
-$\|\partial_{B}^{{\beta}}R^{{\varepsilon}}(B,G)\|\leq C_{{\beta}}M{\varepsilon}^{-|{\beta}|-1}$
-where $M=\sup\{\,\|G(s)\|:|\operatorname{Im}s|\leq{\varepsilon}\,\}$ and the constants are independent of ${0<\varepsilon}\leq 1$.
-Observe that for the general case when $B$ is not symmetric, we find that the estimates (4.5) have the factor $N$ in front of $|\beta|$, see [2, Lemma A.1].
-Proof.
-This will be an adaptation of the proof of Lemma A.1 in [2].
-Let $A$ be symmetric $N\times N$ matrix, and assume $\|A\|<1$.
-This gives that 1<t<1-1<t<1 when $|P(t,A)|=0$ since then tw=Awtw=-Aw for some $w\neq 0$.
-Thus we have that
-(B.6) $|P(t,A)|\neq 0\qquad\text{when $\operatorname{Im}t\neq 0$}$
-Let $\omega$} be the rectangle $\{\,s:|\operatorname{Im}s|<{\varepsilon},|\operatorname{Re}s|<2\,\}$, $0<\varepsilon\leq 1$, and let ${\omega}_{t}$ be the union of this rectangle and a congruent one centered at $\operatorname{Re}t$.
-Then (B.1) is satisfied when $|\operatorname{Im}t|<{\varepsilon}$, with
-(B.7) $Q(t)=(2{\pi}i)^{-1}\int_{\partial{\omega}_{t}}G(s)P(s,A)^{-1}(s-t)^{-1}\,ds$
-and
-(B.8) $R=(2{\pi}i)^{-1}\int_{\partial{\omega}}G(s)P(s,A)^{-1}\,ds$
-Since $P(t,A)$ is symmetric, we find from (B.6)
-(B.9) $\|P(s,A)^{-1}\|\leq{\varepsilon}^{-1}\quad\text{on}\quad\partial{\omega}_{t}\bigcup\partial{\omega}$
-In fact, if $\alpha_{j}\in(-1,1)$ are the eigenvalues of $A$ then the eigenvalues of $P(t,A)$ are $t+\alpha_{j}$, and $\|P(t,A)^{-1}\|\leq\max_{j}|t+\alpha_{j}|^{-1}\leq\varepsilon^{-1}$ when $|\operatorname{Im}t|\geq\varepsilon$ or $|\operatorname{Re}t|\geq 2$.
-This gives the estimates
-(B.10) $\left\{\begin{aligned} &\|Q(t)\|\leq CM{\varepsilon}^{-2}\quad\text{when $|\operatorname{Im}t|<{\varepsilon}$}\\ &\|R\|\leq CM{\varepsilon}^{-1}\end{aligned}\right$.
-Now we shall perturb the coefficients of $P$.
-Since $\|P(s,A)^{-1}\|\leq{\varepsilon}^{-1}$ on $\partial{\omega}\cup\partial{\omega}_{t}$, we find $\|P(s,B)^{-1}\|\leq c{\varepsilon}^{-1}$ there, if $\|B-A\|\leq c_{0}{\varepsilon}$ for some $c_{0}\ll 1$.
-In fact, the eigenvalues of symmetric matrices depends Lipschitz continuously on the matrix norms by the minimax principle.
-Thus, by using (B.7)–(B.8) for P(t,B)P(t,B)\/ we obtain (B.4) with $Q^{{\varepsilon}}=Q$ and $R^{{\varepsilon}}=R$ when $\|B-A\|\leq c_{0}{\varepsilon}$.
-Clearly, QQ\/ and RR\/ are analytic functions of $B$.
-We easily obtain the estimates (B.5) when $|\operatorname{Im}t|<{\varepsilon}$ and $\|B-A\|\leq c_{0}{\varepsilon}$.
-Next, take a suitable partition of unity $\{\,{\Psi}_{j}(X)\,\}$ for the symmetric matrices, which can be parametrized by $\mathbf{R}^{N}\times\mathbf{C}^{N(N-1)/2}\cong\mathbf{R}^{N^{2}}$, for example the one in [5, Theorem 1.4.6], and put ${\Phi}_{j}(X)={\Psi}_{j}(X/c{\varepsilon})$.
-If $c$ is small enough (independently of $\varepsilon$}), we obtain $\|X-Y\|<c_{0}{\varepsilon}$ if $X$, $Y\in\operatorname{\rm supp}{\Phi}_{j}$, $\forall\,j$.
-By doing the above construction with $A_{j}\in\operatorname{\rm supp}{\Phi}_{j}$, we get the functions $Q^{{\varepsilon}}_{j}$ and $R^{{\varepsilon}}_{j}$.
-By taking $Q^{{\varepsilon}}=\sum_{j}{\Phi}_{j}Q^{{\varepsilon}}_{j}$ and $R^{{\varepsilon}}=\sum_{j}{\Phi}_{j}R^{{\varepsilon}}_{j}$, we obtain the result.
-∎
-The proof of Proposition A.2 in [2] now gives the following result.
-Proposition B.2.
-Let $G(t)\in\mathcal{S}(\mathbf{R})$ with values in $N\times N$ matrices and let $B$ be a symmetric $N\times N$ matrix.
-Then we can find smooth $N\times N$ systems $Q(t,B,G)$ and $R(B,G)$ depending linearly on $G(t)$, such that (B.4) holds with QQ\/ and RR\/ instead of $Q^{{\varepsilon}}$ and $R^{{\varepsilon}}$.
-Also, we have the following estimates
-(B.11) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q(t,B,G)\|\leq C_{{\alpha}{\beta}}\int(\|G\|+\|G^{(\nu)}\|)\,dt{\quad}\nu={\alpha}+|{\beta}|+4$
-$\|\partial_{B}^{{\beta}}R(B,G)\|\leq C_{{\beta}}\int(\|G\|+\|G^{(\nu)}\|)\,dt{\quad}\nu=|{\beta}|+3$
-Observe that in the general case, when $B$ is not symmetric, the number of derivatives of $G$ in (B.11) grows as $N|\beta|$ so the estimates are not tame in general, see Appendix C.
-For completeness, we shall prove Proposition B.2.
-Proof.
-Choose a function $\psi\in C_{0}^{\infty}$ such that $0\leq\psi\leq 1$, $\psi$ has support where $|t|\leq 2$ and $\psi=1$ on [1,1][-1,1]. Let
-(B.12) $G_{0}(t)=(2\pi)^{-1}\int\widehat{G}(\tau)\psi(\tau)\,d\tau$
-(B.13) $G_{j}(t)=(2\pi)^{-1}\int\widehat{G}(\tau)\big(\psi(2^{-j}\tau)-\psi(2^{1-j}\tau)\big)\,d\tau\qquad j\in\mathbf{Z}_{+}$
-Clearly, $G_{j}$ is analytic $N\times N$ matrix and that $G=\sum_{j=0}^{\infty}G_{j}$. We also have
-(B.14) $|\tau^{\nu}\widehat{G}(\tau)|\leq\int|G(t)|+|G^{(\nu)}(t)|\,dt$
-which gives
-(B.15) $|G_{j}(t)|\leq C2^{j(1-\nu)}\int|G(s)|+|G^{(\nu)}(s)|\,ds\qquad\text{if $|\operatorname{Im}t|<\varepsilon_{j}=2^{-j}$}$
-Thus we find from Lemma B.1 that
-(B.16) $G_{j}(t)=Q^{{\varepsilon}_{j}}(t,B,G)P(t,B)+R^{{\varepsilon}_{j}}(B,G){\quad}\text{if}{\quad}\|B\|<1$
-where
-(B.17) $\|\partial_{t}^{{\alpha}}\partial_{B}^{{\beta}}Q(t,B,G)\|\leq C_{{\alpha}{\beta}}2^{-j(\nu-{\alpha}-|{\beta}|-3)}\int(\|G\|+\|G^{(\nu)}\|)\,dt$
-$\|\partial_{B}^{{\beta}}R(B,G)\|\leq C_{{\beta}}2^{-j(\nu-|{\beta}|-2)}\int(\|G\|+\|G^{(\nu)}\|)\,dt$
-With $\nu$ chosen as in (B.11) we find by summing up the convergent series, that
-(B.18) $Q(t,B,G)=\sum_{j=0}^{\infty}Q^{{\varepsilon}_{j}}(t,B,G)\text{ and }R(t,B,G)=\sum_{j=0}^{\infty}R^{{\varepsilon}_{j}}(t,B,G)$
-are smooth functions such that (B.4) and (B.11) hold. ∎
-Now, we can prove the following symmetric version of the Malgrange division theorem.
-Theorem B.3.
-Let $G(t,x)\in C^{\infty}_{0}(\overline{\omega})$ with values in $N\times N$ matrices, where $\omega$ is a neighborhood of the origin in $\mathbf{R}\times\mathbf{R}^{n}$, then for any $M(x)\in C^{\infty}(\overline{\omega})$ with values in symmetric $N\times N$ matrices we find
-(B.19) $G(t,x)=Q(t,x)P(t,M(x))+R(x)\qquad(t,x)\in\omega$
-where QQ\/ and RR\/ are smooth $N\times N$ systems depending linearly on $G$ satisfying
-(B.20) $\|\partial_{t}^{{\alpha}}\partial_{x}^{{\beta}}Q(t,x)\|\leq C_{{\alpha}{\beta}}\|G\|_{j}{\quad}j={\alpha}+|{\beta}|+4$
-$\|\partial_{x}^{{\beta}}R(x)\|\leq C_{{\beta}}\|G\|_{k}{\quad}k=|{\beta}|+3$
-Here $\|F\|_{j}=\sum_{|\alpha|+|\beta|\leq j}\|\partial_{t}^{{\alpha}}\partial_{x}^{{\beta}}F\|$ is the $L^{2}$ Sobolev norm.
-Observe that Theorem B.3 gives that the mapping
-$(Q(t,x),R(x))\mapsto G(t,x)=Q(t,x)P(t,M(x))+R(x)$
-is a $C^{\infty}$ surjective tame mapping, having a $C^{\infty}$ tame right inverse of degree 4, see Appendix C, but that mapping is not tame for Malgranges preparation theorem in general.
-Observe that the symmetric scalar case when $N=1$ follows directly from the inverse function theorem.
-Proof.
-If $G(t,x)\in C^{\infty}_{0}(\mathbf{R}\times\mathbf{R}^{n})$, then by using Proposition A.2 we obtain (B.19) with $Q(t,x)=Q(t,M(x),G(\cdot,x))$ and $R(x)=R(M(x),G(\cdot,x))$.
-Since $Q(t,x)$ and $R(x)$ depend linearly on $G(t,x)$ we obtain (B.20) from (B.11) by Leibniz’ rule, since any $L^{1}$ Sobelev norm of test functions is bounded by the corresponding $L^{2}$ norm.
-∎
-In the preparation theorems, we have only done right symmetric division by matrix valued polynomials.
-By taking transposes we also obtain the results for left symmetric divisions, this does not affect the norms of the matrices.
-Corollary B.4.
-Let $G(t,x)\in C^{\infty}_{0}(\overline{\omega})$ with values in $N\times N$ matrices, where $\omega$ is a neighborhood of the origin in $\mathbf{R}\times\mathbf{R}^{n}$, then for any $M(x)\in C^{\infty}(\overline{\omega})$ with values in symmetric $N\times N$ matrices we can write
-(B.21) $G(t,x)=P(t,M(x))Q(t,x)+R(x)\qquad(t,x)\in\omega$
-where QQ\/ and RR\/ are smooth $N\times N$ systems depending linearly on $G$ satisfying (B.20).
-We find that (B.21) gives a $C^{\infty}$ surjective tame map of degree 0 having a $C^{\infty}$ tame right inverse of degree 4, see Example C.13.
-Appendix C Tame Fréchet spaces and the Nash–Moser theorem
+Appendix B Tame Fréchet spaces and the Nash–Moser theorem
 In this section, we will recall the Nash–Moser Theorem and the theory of tame Fréchet spaces and tame maps from Sections II and III in [5].
@@
 But for the Nash–Moser Theorem one needs Fréchet spaces with a more refined structure.
-Definition C.1.
+Definition B.1.
 A grading on a Fréchet space is a collection of seminorms $\{\|\cdot\|_{n},\ n\in\mathbf{N}\}$ that defines the topology and is increasing: $\|f\|_{j}\leq\|f\|_{k}$ when $j\leq k$.
@@
 Closed subspaces of graded spaces are graded with the induced norms, and Cartesian products $F\times G$ can be graded with the norms $\|(f,g)\|_{n}=\|f\|_{n}+\|g\|_{n}$.
-Definition C.2.
+Definition B.2.
 Two gradings $\|\cdot\|_{n}$ and $\|\cdot\|^{\prime}_{n}$ are tamely equivalent of degree $r$ and base $b$ if
-(C.1) $\|f\|_{n}\leq C_{n}\|f\|^{\prime}_{n+r}\quad\text{and}\quad\|f\|^{\prime}_{n}\leq C^{\prime}_{n}\|f\|_{n+r}$
+(B.1) $\|f\|_{n}\leq C_{n}\|f\|^{\prime}_{n+r}\quad\text{and}\quad\|f\|^{\prime}_{n}\leq C^{\prime}_{n}\|f\|_{n+r}$
 for any $n\geq b$. If this holds for any $n$ then they are equivalent of degree $r$.
 For example, if $X$ is a compact manifold, then the $C^{k}$ norms, Hölder norms and Sobolev norms are tamely equivalent.
-Definition C.3.
+Definition B.3.
 A linear map $L:\ F\mapsto G$ of a graded space into another is tame, if it satisfies a tame estimate of degree $r$ and base $b$:
-(C.2) $\|Lf\|_{n}\leq C_{n}\|f\|_{n+r}$
-for any $n\geq b$. If (C.2) holds for any $n$ then $L$ is tame of degree $r$.
-Example C.4.
+(B.2) $\|Lf\|_{n}\leq C_{n}\|f\|_{n+r}$
+for any $n\geq b$. If (B.2) holds for any $n$ then $L$ is tame of degree $r$.
+Example B.4.
 A $C^{\infty}$ linear partial differential operator $L$ of order $m$ on $C^{\infty}$ is a tame of degree $m$ since $\|Lf\|_{n}\leq C_{n}\|f\|_{n+m}$ for the $C^{k}$ norms for any $n$.
@@
 Solutions to invertible linear systems of smooth equations give tame maps of degree 0 by Leibniz’ rule.
-Similarly, for any $C^{\infty}$ system $M(x)$, we find that the mapping
-$(Q(t,x),R(x))\mapsto G(t,x)=Q(t,x)(t\operatorname{Id}_{n}+M(x))+R(x)$
+Similarly, for any $m\in\mathbf{N}$ and $C^{\infty}$ system ${\mathbb{M}}(t,x)$, we find that the mapping
+$(Q(t,x),\mathbb{R}(t,x))\mapsto G(t,x)=Q(t,x)(t^{m}\operatorname{Id}_{n}+{\mathbb{M}}(t,x))+\mathbb{R}(t,x)$
 is tame of degree 0.
-Example C.5.
+Example B.5.
 The linear mapping $L:C^{\infty}[0,1]\mapsto C^{\infty}[-1,1]$ given by $Lf(x)=f(x^{2})$ is tame with $\|Lf\|_{n}\leq C_{n}\|f\|_{n}$.
@@
 Compositions of linear tame maps are linear and tame, and $L$ is a tame isomorphism if $L$ is linear isomorphism such that $L$ and $L^{-1}$ are tame.
-Definition C.6.
+Definition B.6.
 Let $F$ and $G$ be graded spaces.
 Then we say that $F$ is a tame direct summand of $G$ if we can find tame linear maps $L:\ F\mapsto G$ and $M:\ G\mapsto F$ so that the composition ML:FFML:\ $\mapsto F$ is the identity: $F\overset{L}{\mapsto}G\overset{M}{\mapsto}F$.
-Definition C.7.
+Definition B.7.
 We say that a graded space is tame if it is a tame direct summand of a space $\Sigma(B)$ of exponentially decreasing sequences in some Banach space $B$.
 A Cartesian product of two tame spaces is tame.
-Theorem C.8.
+Theorem B.8.
 The space $C^{\infty}(X)$ is tame if $X$ is a compact manifold (with or without boundary), and if $V$ is a vector bundle over $X$ then the sections of the vector bundle $C^{\infty}(X,V)$ is also tame, since it can be written as a tame direct summand of a tame space.
-Theorem C.9.
+Theorem B.9.
 If $X$ is a compact manifold with boundary, then $C^{\infty}_{0}(X)$ and $C^{\infty}_{0}(X,V)$ are tame.
 Next, we shall look at nonlinear maps.
-Definition C.10.
+Definition B.10.
 Let $F$ and $G$ be graded spaces and $P:\ F\supseteq U\mapsto G$ a nonlinear map, then we say that $P$ satisfies a tame estimate of degree $r$ and base $b$ if
-(C.3) $\|P(f)\|_{n}\leq C_{n}(1+\|f\|_{n+r})$
+(B.3) $\|P(f)\|_{n}\leq C_{n}(1+\|f\|_{n+r})$
 for all $f\in U$ and all $n\geq b$.
-If (C.3) holds for any $n$ then $P$ is tame of degree $r$.
+If (B.3) holds for any $n$ then $P$ is tame of degree $r$.
 We say that $P$ is a tame map if $P$ is defined on an open set, is continuous and satisfies a tame estimate in a neighborhood of any point.
-As before, if (C.3) holds for any $n$ then we say $P$ satisfies a tame estimate of degree $r$.
-Theorem C.11.
+As before, if (B.3) holds for any $n$ then we say $P$ satisfies a tame estimate of degree $r$.
+Theorem B.11.
 A composition of tame maps is tame.
-The following is the Nash–Moser theorem that we shall use.
-Theorem C.12.
+The following is the Nash–Moser theorem that we shall use, which is Theorem 1.7.1 in section III of [5].
+Theorem B.12.
 Let $F$ and $G$ be tame spaces and $P:\ F\supseteq U\mapsto G$ a $C^{\infty}$ tame map of the open set $U$.
@@
 If the derivative $DP(f)$ is injective for any $f\in U$ and the family of left inverses $VP(f):\ U\times G\mapsto F$ are tame maps, then $P$ is locally injective and each local left inverse is a $C^{\infty}$ tame map.
-Example C.13.
-It follows from Corollary B.4 that the mapping
-$C^{\infty}(\mathbf{R}\times\mathbf{R}^{n})\times C^{\infty}(\mathbf{R}^{n})\ni(Q(t,x),M(x))\mapsto G(t,x)=P(t,M(x))Q(t,x)+R(x)$
-with $M(x)=M^{*}(x)\ \forall\,x$ is a $C^{\infty}$ tame surjective map of degree 0 with local $C^{\infty}$ tame right inverses of degree 4.
+Remark B.13.
+The local inverse $P^{-1}$ of $P$ in Theorem B.12 inherits the properties of the linear inverse VPVP, since $VP$ is used in the construction of $P^{-1}$, see Theorem 1.7.1 in section III of [5].
+For example, if VPVP depends $C^{\infty}$ on some parameters then $P^{-1}$ also depends $C^{\infty}$ on these.
+Example B.14.
+It follows from Proposition 4.2 that if ${\mathbb{M}}(t)$ is a smooth symmetric $N\times N$ system valued function such that ${\mathbb{M}}(0)=0$ and $\partial_{t}^{m}{\mathbb{M}}\equiv 0$ then the mapping of smooth $N\times N$ system valued functions
+(B.4) $C_{0}^{\infty}(\mathbf{R})\times C_{0}^{\infty}(\mathbf{R})\ni(Q(t),{\mathbb{M}}(t))\mapsto\\ G(t)=P(t,{\mathbb{M}}(t))Q(t)+R(t)\in C_{0}^{\infty}(\mathbf{R})\times C_{0}^{\infty}(\mathbf{R})$
+is a $C^{\infty}$ tame surjective map of degree 0 with a local $C^{\infty}$ tame right inverse of degree $3+m$.
 The proofs of these results and more examples and counterexamples can be found in sections II and III in [5].
@@
 I would like to thank Gregory Berkolaiko at Texas A&M, who asked me about symmetric preparation theorems like (1.5) which started this project.
-We then worked together on the formal power series solution in Section 3.
-I am also indebted to Dmitry Kerner at Ben-Gurion University, who informed us about the nested Artin Approximation, which gives analytic solutions in the two dimensional case from the formal power series solution, see Remark 2.2.
+We then worked on the formal power series solution in the first order case, see [4, Section 2].